### v1

# MAAdvisor: Zero-Shot Index Advisor using Multi-Agent LLMs

Abstract.
Index recommendation is one of the most important problems in database management system (DBMS) optimization. Given queries and certain index-related constraints, traditional methods rely on heuristic optimization or learning-based models to select effective indexes and improve query performance. However, heuristic optimization suffers from high computation time, and learning-based models lose generalisability due to training for different workloads and database schemas. With the recent rapid development of large language models (LLMs), methods using prompt tuning have been proposed to enhance the efficiency of index selection. However, such methods still can not achieve the state-of-the-art (SOTA) results, and preparing the index selection demonstrations is also resource-intensive. To address these issues, we propose MAAdvisor, a zero-shot LLM-based index advisor with a multi-agent framework. We decompose the index recommendation problem into sub-steps, including planning, selection, combination, revision, and reflection. A set of LLM-embedded agents is designed to handle each one of the different sub-steps. Our method utilizes global agents to control the index selection process and local agents to select and revise indexes. Through extensive experiments, we show that our proposed MAAdvisor not only achieves the SOTA performance compared to the heuristic methods, but also outperforms learning-based and prompt-based methods with higher efficiency and better zero-shot inference ability.
1. Introduction
Indexes in database systems significantly accelerate data retrieval and reduce redundant I/O operations by providing alternative access paths to data based on attribute values. With the appropriate index configuration, database systems can better support analytical workloads that are highly demanded by many modern applications, such as inventory management, online shopping, and real-time advertising. However, selecting and creating indexes always comes with significant computation time and storage cost. Therefore, given the constraints such as limited resource and budget, it is important for database administrators (DBAs) to select the optimal indexes that can meet the data management needs.
| Heuristic-based method | Learning-based method | Prompt-based method | MAAdvisor (Ours) | |
| Zero-shot Generalization | Yes | No | No | Yes |
| Effectiveness | High | Medium | Medium | High |
| Efficiency | Low | Low | High | High |
To assist DBAs with index selection, index advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Lan et al., 2020; Perera et al., 2021; Sadri et al., 2020a, b; Skelley, 2000; Wang et al., 2024b; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997; Kossmann et al., 2022; Zhao et al., 2025; Zhou et al., 2022; Paludo Licks et al., 2020) are often adopted to automatically select the optimal indexes for a given query workload on the target database. Existing index advisors fall into three main categories: heuristic-based advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Skelley, 2000; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997), learning-based advisors (Lan et al., 2020; Perera et al., 2021; Sadri et al., 2020a, b; Wang et al., 2024b; Kossmann et al., 2022; Zhou et al., 2022; Paludo Licks et al., 2020), and prompt-based LLM advisors (Zhao et al., 2025). Heuristic-based advisors typically use rule-based or greedy algorithms to identify optimal indexes based on database optimizers’ cost estimations. However, given the large number of columns involved in complex workloads, exploring exhaustively the search space for all possible index combinations becomes prohibitory expensive. In addition, the reliance on heuristics and optimizers’ estimations can lead to suboptimal index selection. As highlighted in (Ding et al., 2019), the errors in cost estimations can even cause workload performance degradation, a phenomenon known as the index regression problem. Learning-based advisors have emerged as a promising alternative to traditional heuristic-based approaches. They utilize techniques like deep learning and reinforcement learning to train index selection models and index benefit estimators. Through extensive learning on features extracted from queries, execution plans, and database statistics, these models can capture complex interactions between indexes and queries that are difficult to encode manually and hence achieve SOTA performance. However, training these models often requires a large amount of labeled data and time. The lack of ability to generalize to unseen database limits their applicability in dynamic or evolving environments. Moreover, the need for data annotations and model re-training poses a major limitation in data-sensitive domains like healthcare or security, where data collection and sharing are tightly regulated. To increase advisors’ generalisation and reduce recommendation time cost, prompt tuning on LLMs is adopted (Zhou et al., 2023). Such a method prompts LLMs with preprocessed index recommendation demonstrations and lets the LLM select columns to build the indexes. Although the generalisation and efficiency of advisors benefit from LLM, there are some limitations. First, the performance of the model is highly sensitive to the prompt design, such as the choice and format of demonstrations. Second, demonstrations on the analyzed database schema are required to achieve optimal results, therefore, preparing high-quality demonstrations is time-consuming and often requires expert domain knowledge. Lastly, inference with LLMs can be computationally expensive for large-scale workloads. Table 1 summarizes the characteristics and drawbacks of three types of advisors as discussed above.
Given the limitations of existing methods, we aim to design an index advisor that addresses the following fundamental challenges in index recommendation: (1) Zero-shot Generalization: To accommodate the diverse landscape of real-world database deployments, an effective index advisor must exhibit strong generalization capabilities. Specifically, the ability to recommend optimal indexes for previously unseen databases and workloads without requiring retraining or fine-tuning on domain-specific data. (2) Effectiveness: The advisor should deliver SOTA recommendation quality, producing index configurations that demonstrably improve query performance compared to existing approaches. (3) Efficiency: The advisor must exhibit low computational overhead, enabling rapid index recommendations that scale to large workloads without imposing significant time costs.
To address the challenges above, we propose our multi-agent index advisor based on LLM, named MAAdvisor. First, to address the zero-shot generalization challenge, we build upon the success of recent multi-agent work solving other problems under the zero-shot setting (Liu et al., 2025; Li et al., 2025; Jha et al., 2025). Multi-agent frameworks can enhance scalability by dividing complex problems into manageable components processed in parallel by specialized zero-shot agents. To address the difficulty of decomposing the index recommendation problem into sub-problems, we innovatively design a hierarchical multi-agent pipeline including global agents and local agents. Global agents decompose the problem into iterative sub-steps and plan the use of local agents in each sub-step. Local agents will then perform three key functions: (1) select candidate index, (2) revise the current index selection, and (3) construct composite indexes. This design exploits LLMs’ demonstrated reasoning capabilities in index recommendation and enables the system to generalize across diverse database schemas and workload patterns in zero-shot. Second, to solve the effectiveness challenge caused by the index regression problem in most advisors, we design two novel components for our index advisor: (1) a new workload representation paradigm for the local selection agent to accurately select optimal indexes, and (2) a dedicated local revision agent that identifies and eliminates potentially regressive index recommendations by analyzing estimation errors and performance trade-offs. Third, to guarantee the computational efficiency, we constrain the search space through a budget-aware planning mechanism where the global planning agent uses storage constraints as an optimization objective, effectively pruning the solution space explored by local agents.
We conduct extensive experiments and recommend indexes for multiple popular datasets and workloads, including TPC-H, TPC-DS, DSB, and JOB. We observe that MAAdvisor’s benefit-to-cost ratio outperforms all the heuristic-based baselines, which are also zero-shot, in 10 out of the 16 settings. In addition, MAAdvisor significantly outperforms the learning-based methods, which require intensive training. Compared with the few-shot prompt-based baseline, MAAdvisor also achieves better performance with lower inference time and no domain-specific data.
Our main contributions are as follows:
To the best of our knowledge, we are the first to propose MAAdvisor, a zero-shot index advisor using LLMs, achieving SOTA results with higher efficiency and zero-shot generalization.
We propose a new workload representation paradigm that is suitable for LLMs to understand and analyze the index recommendation problem.
We design a multi-agent framework to solve the index recommendation problem. The problem is decomposed into multiple sub-steps, which are much easier to solve by a cooperation of the global and local agents.
2. Preliminary
In this section, we will first clarify the concepts related to index recommendation in Section 2.1. Next, we will then explain the problem of an index advisor in Section 2.2.
2.1. Workload, Index Candidate and Budget
Workload: The workload is referred to as a set of queries that will be executed in the database. The ultimate goal of the index recommendation problem is to optimize the execution efficiency of the entire workload instead of any specific query within it.
Index Candidate: Most of the methods that solve the index recommendation problem start with collecting a set of index candidates to select from. Each candidate within this set usually stands for a table column queried by any of the queries in the workload.
Storage Budget: Index selection in databases is subject to various practical constraints, most notably storage budget. Creating indexes consumes additional disk space and incurs extra cost during data modifications. During index selection, it is significant to balance performance gains against the increased resource usage, which necessitates prioritizing indexes that offer the highest benefit-to-cost ratio.
2.2. Index Advisor
Index advisors receive the workload of a specified database along with constraints like the maximum number of allowable indexes and storage budget, and then automatically generate an index recommendation. The functionality of an index advisor can be formally defined as follows:
Definition 2.1.
(Index Advisor): Given a database , a workload consisting of queries querying the database and a storage budget , an index advisor will automatically compute the column candidates and construct a set of indexes with columns , optimizing the cost of the workload by
where calculates the execution time of the workload querying the database with the indexes created, and calculates the total storage cost to construct the indexes selected.
3. Related Work
3.1. Index Advisors
Heuristic-based Index Advisors: Heuristic-based advisors iteratively and greedily select candidate indexes until the budget constraint is met. Within each optimization step, the index advisors will update the index recommendation according to the estimated benefits of the index candidates. There are mainly two heuristics for the index advisors to follow: the ADD heuristic and the DROP heuristic. As an example of the ADD heuristic, Extend (Schlosser et al., 2019) starts with no indexes and adds one index at a time according to the highest benefit-to-cost ratio. In contrast, Drop (Choenni et al., 1993), following the DROP heuristic, starts with all index candidates and iteratively removes the least beneficial one until optimization constraints are met. In general, heuristic-based advisors generally take significantly longer computation time to recommend indexes, especially for databases with complicated schemas. In addition, heuristic-based methods rely solely on database optimizers to evaluate index benefits. According to (Ding et al., 2019), wrong estimations in column cardinality may mislead index advisors to select indexes that are not beneficial to the workload.
Learning-based Index Advisors: While recent work (Zhou et al., 2024) shows that supervised learning approaches such as QueryFormer (Zhao et al., 2022) can be applied to index recommendation, state-of-the-art learning-based advisors (Kossmann et al., 2022; Wang et al., 2024b; Lan et al., 2020; Sadri et al., 2020a; Paludo Licks et al., 2020) predominantly adopt reinforcement learning (RL) algorithms. For example, DQN (Lan et al., 2020) integrates Q-learning with deep neural networks to derive optimal policies from workload representations, while SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b) leverage the PPO (Schulman et al., 2017) algorithm for policy optimization, with BALANCE further employing a transfer learning framework to exploit historical workloads. RL-based advisors typically achieve much lower inference latency than heuristic-based counterparts; however, the training time for such methods cannot be overlooked. In addition, the generalisability of the RL-based methods is lower since training is required for each unseen database.
Prompt-based Index Advisors: Large Language Models (LLMs) have recently emerged as a hot topic in machine learning research. These models have demonstrated a surprisingly strong ability to handle a variety of text-related tasks such as generation, decision-making, and reasoning. Therefore, (Zhao et al., 2025) proposes to utilize LLMs for index recommendation. They construct and offer in-context demonstrations for LLMs to learn from. LLMs will then directly recommend indexes for both met and unmet databases based on the demonstrations. When a new workload is given, LLMIdxAdvis explores the prepared demonstration pool and selects the optimal demonstration for the workload by cosine similarity. LLMIdxAdvis successfully addresses the problem of efficiency and robustness in index recommendation. However, the efficiency and generalisability are achieved at the cost of the effectiveness of the indexes recommended. LLMIdxAdvis achieves only comparable performance against other methods in simpler workloads and underperforms the SOTA methods given more complicated workloads and database schemas.
3.2. Multi-agent Pipelines
Leveraging the strong generalizing and reasoning abilities of LLMs, recent research proposes the idea of agents (Shen et al., 2023; He et al., 2016). LLM agents employ a more complex, modular architecture, enabling end-to-end problem-solving and the ability to handle multi-step tasks. In a standard multi-agent pipeline (Liu et al., 2025; Li et al., 2025; Jha et al., 2025), a complex problem is divided into different sub-problems. The sub-problems are assigned to specialized agents that collaborate in parallel. Each agent will focus on one task, such as data pre-processing, information retrieval, reasoning, or generation. Communication occurs through shared memory or message passing among different agents. The agents can be either rule-based or learning-based, which improves scalability and interpretability, enabling efficient problem-solving for complex workflows requiring coordination, specialization, and iterative refinement among agents.
4. MAAdvisor Pipeline Overview
Figure 1 illustrates the pipeline of MAAdvisor. In this section, we will explain our motivations in the method design and offer an overview of the two key sections in our method, named workload representation (4.1) and multi-agent pipeline (4.2).
| Execution time (s) | With regressions | Remove regressions |
| No index | 193.58 | 193.58 |
| Extend | 161.28 | 160.53 |
| GPT-4o | 186.70 | 155.54 |
MAAdvisor
4.1. Workload Representation
Motivation: To enable zero-shot LLM reasoning effectively and efficiently, two main challenges lie within the information for index recommendation, namely long input context and lack of domain knowledge. First, the database information, like schema, workloads, and query plans, can be extremely long in real-world applications. For example, a workload consisting of 83 queries from the TPC-DS benchmark alone can take 36518 tokens’ length in GPT-4’s input. Not only do smaller LLMs may not have the ability to handle such long input sequences, but the cost and hallucination (Ji et al., 2023; Zhang et al., 2023) for larger LLMs dealing with long contexts can also be considerably large. Therefore, a new representation is required to summarize the database-related information into effective and precise LLM inputs. Second, although previous work (Wei et al., 2023) has proven LLMs’ strong generalization and reasoning ability given inputs similar or related to their training data, general-purpose LLMs always have insufficient training data related to data management and database indexing. Therefore, effective workload representation is vital for solving the index recommendation problem with LLMs.
Unfortunately, the workload representation techniques from existing methods are not fit for LLMs. The learning-based methods usually encode the database and workload information into embedded features (Lan et al., 2020; Wang et al., 2024b). However, such features are too abstract for LLMs to understand and use as inputs. The prompt-based method (Zhao et al., 2025) also attempts to address this problem by extracting index-related information in the workload for the LLMs. However, the information is yet stored in densely structured nested dictionaries. It is difficult for LLMs to locate the information needed, and hallucination may be triggered according to our experiments in Section 7.2. Therefore, a workload representation technique that is able to both shorten the LLM input length and preserve the critical information is required for LLMs to perform the index recommendation task.
Overview: To address the challenges above, we reformulate and summarize the inputs in a textual way that LLMs can easily understand and use. The information from workload, database schema, and optimiser that is relevant to index recommendation is extracted, merged, and grouped by the indexable column candidates. Given all the column candidates and their relevant information affecting index recommendation, the problem of using LLMs to construct indexes from raw workload can be simplified into selecting and combining columns from existing candidate choices. At the same time, by merging inputs on a per-column basis, the LLM’s prompt length is capped at the database’s total column count. In other words, regardless of how many queries appear in the workload, the set of candidate columns can never surpass the number of columns defined in the schema.
4.2. Multi-agent Pipeline
Motivation: To explore the capability of LLMs to solve the index recommendation task, we first conduct a pioneering experiment on the TPC-DS benchmark. We generate 100 small workloads, with each consisting of 3 to 10 random queries querying the TPC-DS database, and set a 2 GB storage budget. The results of directly letting GPT-4o (Achiam and et al., 2023) select indexes in a zero-shot manner given column candidates, and the heuristic-based baseline Extend (Schlosser et al., 2019) are presented in Table 2. It is observed that GPT-4o underperforms Extend by 25.42 seconds on average. Although introducing GPT-4o increases the scalability of the index recommendation, it demonstrates that the problem is yet too complicated for a single call of LLM to solve in zero-shot. Similar findings have been reported in (Zhao et al., 2025) that even with one-shot demonstration, LLMs still fail to outperform existing methods. However, when we remove the regressing indexes from the workloads and only keep the improving ones, it is observed that GPT-4o can outperform Extend. It demonstrates LLM’s potential in index recommendation and also justifies that an efficient and well-designed framework powered by LLM can successfully address the problem. Pioneer experiments also reveal that GPT-4o’s index recommendations fluctuate considerably, highlighting the need for a budget-aware planning pipeline to ensure the LLM produces consistent outputs while adhering to budget constraints.
Overview: To address the challenges above, we propose to simplify the index recommendation problem using a hierarchical multi-agent pipeline with budget-aware planning. At the global level, the multi-agent pipeline reflectively decomposes the problem into LLM-solvable sub-steps, including selecting, revising, or combining column candidates. At the local level, each sub-step is handled by a specialized local agent embedded with LLM. According to language model theory (Brown et al., 2020), LLMs produce more accurate and consistent outputs when handling simpler tasks. By breaking down the complex problem into smaller, well-defined sub-tasks and delegating them to specialized LLM agents, our multi-agent pipeline achieves both stability and efficiency in index recommendation. Moreover, by estimating the storage consumption at each sub-step, subsequent steps are planned accordingly, enabling budget-aware index recommendation.
5. Workload Representation
In this section, we will introduce our workload representation technique. In Section 5.1, we will first introduce the information we extract as a representation for each query in the workload. Then in Section 5.2, we will explain how we reformulate and merge the information in the whole workload into column candidates.
5.1. Query Information for Index Selection
To fully leverage the reasoning capabilities of LLMs, the information input should be concise and readily interpretable. However, the data extracted directly from SQL queries is often highly structured and domain-specific. Therefore, it is necessary to extract and reformulate the query-related information for the LLMs to interpret and process effectively. Building on insights from prior studies (Zhao et al., 2023, 2025; Siddiqui et al., 2022) and our preliminary experiments, we summarize key principles for extracting query-level information that are most relevant and effective for index recommendation with LLM:
Principle 1: To perform zero-shot inference, information that varies with databases and does not significantly affect index selections, for example, the data values in join conditions, predicates, and database knobs, can be excluded from the input;
Principle 2: Only columns that appear in any of the workload queries need to be considered as candidates to construct indexes, which will be called ‘indexable columns’.
Based on the principles above, the information we extract from a query representing each indexable column is as follows:
Table and column name: Table and column names are the key attributes for indexes. In addition, the semantics of the table and column names can usually offer the LLMs additional information during index recommendation. For simplification, the format ‘TableName.ColumnName’ will be used in the column candidate for LLMs to choose during the selection phase;
Column operator type: Operators play important roles in index recommendation. Indexes can post a significant change in the efficiency of an operation. The operators related to the indexable column can be easily accessed in a query plan;
Estimated column cardinality: As one of the key attributes affecting index selection, cardinality represents the uniqueness of data values contained in a column. Generally, indexes on higher cardinality columns allow easier access to specific rows. However, indexes on low cardinality columns may also help in specific situations, for example, the column value queried is extremely infrequent compared to the other values in the same column;
Column distribution: Column distribution includes data histogram, minimum, and maximum values within the column. Combining the column distribution and column cardinality, LLM can have better discrimination of more and less effective indexes;
Estimated column storage: To optimize the efficiency of the indexes selected, it is necessary to keep track of the storage cost of the current index recommendation. Although the real storage cost can not be calculated before the actual index creation, the database optimizers can estimate the storage using hypothetical indexes (hyp, [n.d.]);
Estimated column utility: Apart from the information above, an estimation of the potential improvement from a column index is vital for telling if a column candidate is worth selecting (Siddiqui et al., 2022). We define the utility of a potential index on a query as follows:
Definition 5.1.
(Index Utility): Given an index to create and a query , the utility of creating the index for is defined by query total cost without and with index per unit of storage cost:
where the database optimizer uses to estimate the total time cost of the query , and uses to estimate the storage cost of the index .
The utility of a column on query will then be calculated as . From the calculation, the estimated column utility is designed to be compound information, including column cardinality, query selectivity, and operations on the column.
5.2. Column Candidate Generation
In Section 5.1, we outlined the information required from each query for the index recommendation task. However, there still exist considerable duplications in information derived from each column. For a workload with indexable columns, where each column appears in queries on average, a total of pieces of information are collected. To enable the LLM to provide effective selections, this information must be consolidated so that each candidate item corresponds to a distinct column. It is observed that, among the information collected for a column, only the column operator type and the estimated column utility vary across queries. The column operator type can easily be merged by recording all the operators in a list. Consequently, an algorithm merging the estimated column utility for all queries in the workload can effectively eliminate duplicate column candidates.
Since the goal of index recommendation is to improve workload performance, the index that reduces the execution cost more should be considered as more beneficial. Therefore, we merge the estimated column utility for the same column in different queries by aggregating their utility values. If a column has a higher utility, it means that selecting the column as an index will result in an estimated higher workload performance gain. For the same workload above, there will be exactly column candidates left for the indexable columns after the merging. With the utility merged by the columns, each column candidate is represented as a dictionary of information so that the LLMs can easily access and read. It is also important to note that the merging also benefits the LLM efficiency, especially for larger workloads, by largely reducing the input length.
agents
6. Multi-agent Pipeline
In this section, we will introduce our multi-agent pipeline in which the index recommendation problem is decomposed into sub-steps for agents to solve. As shown in Figure 2, the pipeline consists of two types of agents: local agents and global agents. In Section 6.1, we will first introduce the three local agents that will select, combine, or revise the index recommendation in each sub-step. We will then explain our design for the two global agents planning each sub-step and reflecting on the plans in Section 6.2. In Section 6.3, we will illustrate the whole multi-agent pipeline on how we let the agents cooperate and communicate to perform index recommendation.
6.1. Local Agents
Selection agent: As the key function of our index advisor, the selection agent plays the role of selecting new columns from the column candidates given. The inputs for the selection agent include previous selected indexes, column candidates to select from, total storage budget, and the storage budget already used. In each single call of the selection agent, LLM is prompted with the inputs to select one new column into the current index recommendation as a single-column index.
Combination Agent: As noted earlier, the selection agent produces only single-column indexes. However, when multiple columns from the same table are frequently queried together, a composite index on those columns may deliver better performance than maintaining them as separate single-column indexes. To address this, we design a combination agent that can merge such indexes. To identify optimal composite indexes, we follow the approach in (Siddiqui et al., 2022), collecting potential rules for combining single-column indexes. As illustrated in Figure 3(b), these rules specify possible column combinations and their ordering based on query operations, which can be readily extracted from the queries themselves. In addition, we incorporate external knowledge from diverse sources such as the Internet, expert practices, and textbooks to fully leverage the LLM’s reasoning capabilities, thereby infusing domain expertise and facilitating effective index combination. Examples of external knowledge can also be found in Figure 3(c). Given the input of current selected indexes with operations, index combination rules, and external knowledge, the combination agent will decide if some indexes can be combined into a composite index and output the new index recommendation.
prompt examples
Revision Agent: The revision agent will be adopted to eliminate the potential duplicates incurred and index regressions caused by the current index recommendation. The duplicates can be easily removed by simple coding techniques. However, as addressed by the pioneer experiments in Table 2, index regression is one of the unsolved problems for both traditional and LLM-based index advisors. One of the reasons for the problem is the database optimizer’s estimation error for cardinality, which is used for evaluating the index’s potential benefits. Apart from the optimizer estimations, LLMs can also obtain cardinality-related information from the name and distribution of the columns. For example, the cardinality for the column ‘country_name’ is highly likely to be smaller than the column ‘order_id’. Therefore, we prompt the LLM to identify and remove columns from the selected index when a significant discrepancy exists between the estimated cardinality and the LLM’s analysis. To enhance the reasoning, we incorporate expert knowledge drawn from diverse sources such as common practices and online discussions. Insights of inefficient indexes are then distilled into expert experience rules, which guide the detection of regressive indexes. An example of such rules is shown in Figure 3(d).
To strengthen the LLM’s capacity for independent and robust reasoning beyond reliance on the expert experiences, we also adopt a regression indicator to provide the LLM with an objective, data-driven measure that complements LLM reasoning.
Regression indicator: Prior research (Ding et al., 2019; Shi et al., 2022) has attempted to address the index regression prediction problem using deep learning techniques. They label query plan pairs before and after index creation as either ‘improvement’ or ‘regression’, and train models on these labeled examples to forecast whether a given index will degrade query performance. However, these methods have limitations in handling zero-shot index recommendations at the workload level. Specifically, they focus only on query-level regressions rather than considering the impact on the entire workload. Moreover, their models require retraining for each new dataset, resulting in limited robustness and generalizability.
To address the aforementioned limitations, we devise a workload-level regression indicator that is trained on features invariant to specific databases and workloads. The overall architecture is depicted in Figure 4. The workload’s query plans, both with and without the indexes, are extracted as inputs for the indicator. Each operator in a plan is represented as a node, forming a binary tree for the entire query. Following the approach in (Zhao et al., 2022), we encode each query tree using tree-biased attention, where a node attends only to those within its sub-tree. The root node then serves as the vector representation of the entire query. The workload representation is constructed by concatenating the root vectors of all queries. Finally, given the pair of workload vectors before and after index creation, we employ an MLP classifier to predict potential index regressions from -1 to 1 based on the difference between the two representations. Each index will be assigned a predicted regressive score that the LLM can use as a reference.
Inspired by (Zhao et al., 2023), we extract only the essential features from each tree node that influence index performance, including the operator type, table, column, estimated cardinality, and cost values. Since operators and estimations are inherently database-invariant, they are encoded using one-hot representations and normalization. To enhance the robustness of the regression indicator, we employ Sentence Transformers (Reimers and Gurevych, 2019) to generate database-invariant embeddings for the table and column information.
regression_indicator
6.2. Global Agents
Planning agent: The planning agent plays an important role in the multi-agent pipeline for managing problem sub-steps and determining which local agents to deploy at each step. Given the column candidates, current indexes, total storage budget, and budget already used, the planning agent first evaluates the stop criteria. If these criteria are not met, it outputs an action for the current sub-step: ’Selection’, ’Combination’, or ’Revision’, each invoking a corresponding local agent. When the planning agent determines that the index selection is optimized or that no additional columns can be chosen within the budget, it triggers the ’Stop’ action to conclude the optimization.
Reflection agent: In practical use, it has been observed that planning agent’s strategies for local agents can be inefficient at times, as it may suffer from hallucinations or become trapped in local optima. For instance, the planning agent may repeatedly revise the current index recommendation, invoke the combination agent when no viable combinations exist, or produce invalid actions due to hallucinations. To enhance the planning process, a reflection agent is introduced to evaluate the current path, to identify and correct errors. This helps the planning agent escape local optima and supports more efficient sub-step planning. Given the sequence of previous sub-step actions and the utility of current indexes, the reflection agent will offer suggestions for the next step’s planning. Example of possible suggestions can also be found in Figure 3(a).
6.3. Multi-Agent Pipeline
Once the global and local agents are developed, they can be integrated into a hierarchical multi-agent pipeline, enabling zero-shot scalable index recommendation. As shown in Algorithm 1, the multi-agent pipeline receives the storage budget and the column candidates as inputs. To balance efficiency with effectiveness, we introduce a hyperparameter , which defines the maximum number of planning sub-steps. At the beginning of each sub-step, the planning agent determines the next action. If additional indexes can be selected, the selection agent chooses a single-column index from the effectively represented column candidates. When combining existing indexes into a composite index could further improve workload performance, the combination agent refines the current recommendation by composing these indexes. Since the selected indexes may contain duplicates, regressions, or low-impact choices, the revision agent ensures effectiveness by removing such cases. After each action, the reflection agent reviews prior decisions and guides subsequent planning. Through the planning agent’s reasoning, if the index configuration is deemed optimal, iterations terminate early, and the final selection is output as the index recommendation.
7. Experiments
In this section, we will evaluate our multi-agent index advisor’s performance in terms of effectiveness, efficiency, and zero-shot generalisation ability.
7.1. Experimental Setup
7.1.1. Dataset.
We select four different datasets from both hypothetical and real-world settings for our evaluations.
TPC-H: The TPC-H dataset is a widely used benchmarking dataset for measuring the performance of database management systems. We generate the data with a scale of 10 (around 10 GB). The TPC-H benchmark has in total of 22 different query templates. In order to simplify the analysis and make a fair comparison of the methods, we follow previous methods (Siddiqui et al., 2022; Schlosser et al., 2019) to remove the query templates 2, 17, 20 that contribute significantly to the total execution time of the workload or have little potential improvement from indexes. The test workload for TPC-H has in total of 19 queries randomly generated from the 19 query templates left.
TPC-DS: The TPC-DS dataset is another widely used benchmark with a more complicated database schema and query templates. Data with a scale of 10 (around 10 GB) is generated. The TPC-DS benchmark has 99 different query templates, and for the same reason in the TPC-H benchmark, we remove the query templates 4, 6, 9, 10, 11, 14, 23, 24, 32, 35, 36, 39, 41, 70, 86, and 95. The test workload for TPC-DS consists of 83 queries generated from the 83 query templates left.
DSB: The DSB benchmark (Ding et al., 2021) is derived from the TPC-DS benchmark, while enhancing the TPC-DS with complex data distribution and challenging yet semantically meaningful query templates. DSB has in total of 15 templates for single block queries, 22 templates for multi-block queries, and another 15 templates for SPJ queries, especially evaluating techniques with limited capabilities. In our experiments, we only utilized the general-purpose 37 query templates in the first two groups and constructed a workload of 37 queries randomly generated from the templates.
IMDB (JOB workload): The IMDB dataset (Maas et al., 2011) is a collection of information about movies, TV shows, and actors. Join order benchmark (JOB (Leis et al., 2015)) is an OLAP benchmark querying the IMDB dataset and designed to assess the capabilities of a database management system (DBMS) in optimizing and executing complex real-world join queries. The JOB workload has 113 real-world queries.
7.1.2. Experimental Environments.
We choose the popular DBMS PostgreSQL 14.2 to evaluate all the queries (Pos, [n.d.]). We follow the official documentaries and websites (TPC, [n.d.]; Leis et al., 2015; dsb, [n.d.]) to create PostgreSQL databases for all the above datasets. All the experiments are conducted on an isolated server with two Intel(R) Xeon(R) Platinum 8469C CPUs featuring 46 cores and 96 threads, along with 460GB of RAM.
7.1.3. Baseline Methods.
Heuristic-based baselines: Heuristic-based index advisors can be broadly categorized into two types by their optimization approaches. Those following an ADD strategy incrementally build a set of indexes by selecting candidates one by one. In our experiments, we include Anytime (Chaudhuri and Narasayya, 2020), DB2Advis (Skelley, 2000), Extend (Schlosser et al., 2019), and Relaxation (Bruno and Chaudhuri, 2005) as representatives of this group. On the other hand, AutoAdmin (Chaudhuri and Narasayya, 1997) and Drop (Choenni et al., 1993) serve as baselines for the DROP strategy, where advisors start with an initial set of indexes and iteratively eliminate those deemed inefficient. Learning-based baselines: We adopt two RL-based index advisors as our baselines, namely SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b). While SWIRL includes vast amounts of workloads during training to largely shorten the inference time, BALANCE utilizes transfer RL to increase the generalisability to the dynamic workload scenarios. Prompt-based baselines: The recent work LLMIdxAdvis (Zhao et al., 2025) proposes to recommend indexes by prompting the LLM with prepared demonstrations. Selecting the best demonstration for the input workload and calling the LLM helps LLMIdxAdvis achieve higher efficiency and generalisability. Due to the high computation cost for preparing new demonstrations for our workloads, we directly use LLMIdxAdvis’ prepared demonstrations since they cover all the databases in our experiment.
7.1.4. Evaluation metrics.
We adopt two metrics for our evaluations. The benefit-to-cost ratio metric is first introduced by (Skelley, 2000) and evaluates the absolute workload execution time reduction per unit storage used. The larger the benefit-to-cost ratio, the better an index advisor performs in both workload cost reduction and efficient storage use. The relative workload improvement is defined as the percentage of the workload’s actual execution time decrease before and after the construction of the recommended indexes. We measure all the workload execution time in seconds and storage cost in MegaBytes (MB).
7.1.5. LLM and Training Setting.
We leverage the GPT-4o-0806 (Achiam and et al., 2023) version within the OpenAI API as the default LLM backbone for the agents. Furthermore, we also assess our system’s generalizability across other LLMs, as detailed in Section 8.3. For the index indicator, we train the model on a server with one Nvidia A100 GPU.
| Dataset | TPC-H | TPC-DS | DSB | JOB | ||||||||||||
| Storage Budget | 2% | 10% | 20% | 50% | 2% | 10% | 20% | 50% | 2% | 10% | 20% | 50% | 2% | 10% | 20% | 50% |
| Anytime | 0.089 | -0.010 | 0.415 | 0.198 | 15.285 | 4.588 | 1.494 | 0.821 | 0.063 | 0.040 | 0.019 | 0.147 | 0.979 | -0.297 | 0.095 | 0.080 |
| AutoAdmin∗ | - | 0.714 | 0.714 | 0.202 | 0.746 | 0.768 | 0.080 | 0.047 | 0.130 | 0.022 | 0.017 | 0.006 | - | - | 0.129 | 0.066 |
| DB2Advis | 0.007 | -0.034 | -0.012 | 0.144 | 7.342 | 2.961 | 1.364 | 0.784 | 0.092 | 0.054 | 0.573 | 0.167 | 1.007 | 0.227 | 0.162 | 0.060 |
| Drop∗ | - | 0.716 | 0.715 | 0.154 | 0.749 | 0.767 | 0.102 | 0.117 | 0.201 | 0.021 | 0.008 | 0.103 | - | - | 0.132 | 0.088 |
| Extend | -0.013 | -0.029 | 0.466 | 0.226 | 2.136 | 2.257 | 0.864 | 0.702 | 0.061 | 0.044 | 0.013 | 0.169 | 1.111 | 0.245 | 0.099 | 0.067 |
| Relaxation | 0.080 | 0.007 | 0.444 | 0.218 | 10.141 | 2.561 | 1.969 | 0.709 | 0.068 | 0.041 | 0.012 | 0.149 | 1.378 | 0.240 | 0.086 | 0.071 |
| SWIRL | 0.010 | 0.027 | 0.496 | 0.172 | 0.629 | -0.557 | 0.496 | 0.153 | 0.059 | 0.022 | 0.211 | 0.007 | 0.279 | 0.097 | 0.113 | 0.080 |
| BALANCE | 0.023 | 0.045 | 0.496 | 0.158 | 4.035 | 0.097 | 0.732 | 0.222 | 0.043 | 0.041 | 0.207 | 0.170 | 0.273 | 0.114 | 0.110 | 0.077 |
| LLMIdxAdvis | -0.013 | -0.025 | 0.684 | 0.203 | 2.136 | 0.754 | 1.642 | 0.394 | 0.061 | -0.630 | 0.246 | 0.250 | 0.943 | 0.202 | 0.126 | 0.090 |
| MAAdivsor | 0.220 | 0.587 | 0.587 | 0.299 | 9.726 | 4.756 | 1.594 | 1.594 | 0.919 | 1.224 | 0.598 | 0.344 | 0.617 | 0.452 | 0.125 | 0.094 |
results_figure
7.2. Effectiveness and Generalisation Evaluation
As shown in Table 3, we compare our proposed MAAdvisor’s benefit-to-cost ratio with the baselines on four benchmarks. All the methods are evaluated with four different storage budgets, taking 2%, 10%, 20%, and 50% of the database size. Our result reveals several key observations:
(1) MAAdvisor outperforms all the baselines in most of the situations. The benefit-to-cost ratio for our method is the highest in 10 out of the 16 experiment settings, and remains the top tier for the rest 6 settings. Specifically, MAAdvisor doubles the index scores from the SOTA performance given 2%, 10%, and 50% storage budgets in the most complicated DSB benchmark. In addition, similar improvement can be observed in the 10% and 50% budget settings for the real-world benchmark JOB. It demonstrates that MAAdvisor can generalize to a variety of database schemas and workloads while maintaining excellent performance.
| Dataset | TPC-H | TPC-DS | DSB | JOB | ||||||||||||
| Storage Budget | 2% | 10% | 20% | 50% | 2% | 10% | 20% | 50% | 2% | 10% | 20% | 50% | 2% | 10% | 20% | 50% |
| Anytime | 17.07 | 24.85 | 42.26 | 84.49 | 1846.86 | 1853.05 | 1850.87 | 1861.07 | 1741.27 | 1874.02 | 1865.24 | 1876.40 | 1850.46 | 1837.55 | 2000.60 | 2618.21 |
| AutoAdmin | - | 0.47 | 3.41 | 20.01 | 12.42 | 50.47 | 50.47 | 168.11 | 10.05 | 73.84 | 73.84 | 135.90 | - | - | 61.93 | 1839.31 |
| DB2Advis | 14.31 | 23.81 | 23.36 | 23.49 | 46.00 | 44.63 | 45.58 | 47.71 | 46.81 | 48.21 | 50.22 | 47.80 | 40.58 | 33.72 | 33.88 | 36.54 |
| Drop | - | 12.00 | 26.57 | 80.71 | 2243.51 | 2241.48 | 2241.48 | 2229.48 | 2026.49 | 2229.92 | 2229.92 | 2229.48 | - | - | 1869.58 | 1849.99 |
| Extend | 16.43 | 17.12 | 14.65 | 23.48 | 57.58 | 119.71 | 192.50 | 359.81 | 55.92 | 96.41 | 136.26 | 309.31 | 91.68 | 362.16 | 693.30 | 1520.86 |
| Relaxation | 40.66 | 39.44 | 43.39 | 36.10 | 4987.62 | 4917.65 | 4536.28 | 4356.80 | 1469.11 | 1476.62 | 1385.13 | 1285.31 | 6952.19 | 6708.86 | 6538.98 | 5846.26 |
| SWIRL∗ | 1.09 | 1.30 | 2.44 | 2.42 | 3.73 | 3.35 | 3.68 | 3.71 | 2.59 | 1.88 | 2.78 | 2.06 | 4.69 | 2.95 | 4.56 | 4.26 |
| BALANCE∗ | 1.03 | 1.36 | 2.34 | 2.28 | 3.40 | 3.31 | 3.61 | 3.44 | 2.47 | 1.96 | 2.86 | 2.15 | 4.45 | 2.86 | 4.77 | 4.31 |
| LLMIdxAdvis | 210.23 | 148.95 | 216.63 | 203.83 | 153.35 | 167.76 | 412.48 | 138.95 | 134.77 | 123.45 | 146.10 | 177.62 | 141.86 | 165.49 | 170.04 | 174.96 |
| MAAdivsor | 31.85 | 44.31 | 122.21 | 156.82 | 51.00 | 85.75 | 106.71 | 133.21 | 43.23 | 83.67 | 115.38 | 178.82 | 69.22 | 66.18 | 106.62 | 211.56 |
(2) Index regressions happen for the zero-shot heuristic-based methods, for example, Anytime in TPC-H and DSB under 10% budget. Similar results can be observed for Extend in TPC-H and for SWIRL in TPC-DS under 10% budget. Since heuristic-based methods highly rely on the benefit estimations from optimizers, index regressions always happen when the index advisors have a limited storage budget and select the regressive indexes. As an example, DB2Advis recommends indexes for TPC-H that lead to a nearly 50% execution time regression on query template 9 from 14.76 seconds to 26.60 seconds. Among the 13 indexable columns in query 9, only two columns appear in DB2Advis’ index recommendation. In addition, the composite indexes DB2Advis selects ‘I(C supplier.s_nationkey,C supplier.s_suppkey)’ and ‘I(C part.p_type,C part.p_partkey)’ are not efficient for JOIN conditions ‘ps_suppkey = l_suppkey’ and ‘p_partkey = l_partkey’. Unfortunately, the improvements for the indexes on other queries are not significant and therefore result in an overall regression. In contrast, MAAdvisor successfully alleviates the index regression problem while maintaining the zero-shot generalisation ability, with the help of the specifically designed workload representation technique and local revision agent.
(3) Aligned with (Zhou et al., 2024), it is observed that the performance for the learning-based index advisors suffers from higher fluctuation across benchmarks compared to heuristic-based methods. As an example, while BALANCE performs well on TPC-DS and DSB, it fails to beat heuristic-based baselines for the rest of the benchmarks. The main reason is that the training process for the learning-based method is highly variable. The hallucination of LLMs is also a similar issue that may affect the variance of the performance. As an example, LLMIdxAdvis’ performance on TPC-DS 20% budget is among the top tier, but the performance on the other budget settings is low. In contrast, with the help of the hierarchical multi-agent pipeline, our zero-shot MAAdvisor maintains a higher performance across the datasets without training and is robust against hallucination.
In order to further analyze the capability of our method, we plot the two metrics we used to compute the benefit-to-cost ratio, namely the relative workload improvement and the storage cost. From Figure 5, we can observe that:
| Avg. Tokens (Cost $) | TPC-H | TPC-DS | DSB | JOB |
| MAAdvisor | 12165(0.03) | 68035(0.17) | 66926(0.17) | 33117(0.08) |
| LLMIdxAdvis | 38736(0.10) | 38859(0.10) | 37361(0.10) | 35619(0.09) |
(1) MAAdvisor achieves the highest relative workload improvement in 8 out of in total 16 experiment settings. Given smaller storage budgets from 2% to 20%, MAAdvisor tends to recommend indexes that lead to the most workload relative reduction. However, for a larger budget like 50%. Our method prefers indexes with comparable improvement but much lower storage cost. It is also observed that in the experiment settings where our method did not achieve the highest benefit-to-cost ratio, for example, TPC-H 20% budget and JOB 20% budget, MAAdvisor can still recommend indexes that result in the largest workload execution time decrease, with a trade-off of higher storage cost.
(2) Heuristic-based baselines following the DROP heuristic fail to perform well for small storage budgets. The main reason is that the small-storage indexes are generally not as beneficial as large-storage indexes and are removed from the recommendation in early DROP stages. Our MAAdvisor assembles the ADD heuristic that iteratively selects indexes into the recommendation by calling the selection agent. Therefore, MAAdvisor can achieve SOTA performance even for small storage budgets.
(3) The prompt-based method LLMIdxAdvis performs well on TPC-H and JOB benchmarks but has a lower performance on TPC-DS and DSB. As also specified in (Zhao et al., 2025), directly calling LLMs to select indexes fails to excel in these complicated workloads and database schemas. In contrast, MAAdvisor outperforms all the baselines, especially in DSB, with our special design. The workload representation technique efficiently summarizes the information from workloads and database schemas. In addition, the multi-agent pipeline decomposes the problem into sub-steps handled by specialized agents, simplifying the task for each LLM agent to solve.
(4) We find LLMIdxAdvis also struggles to control the storage budget used in TPC-DS and DSB. In TPC-DS, LLMIdxAdvis uses 1356.38 MB under 10% budget, which exceeds the budget, and a lower 1122.97 MB under 20% budget. The main reason is that although the available storage budget is an input for LLM in LLMIdxAdvis, the inputs do not specify the storage cost for selecting new indexes. In contrast, MAAdvisor’s planning and selection of indexes is always storage-aware since the estimated storage cost is recorded in each one of the column candidates.
queryid
7.3. Efficiency Evaluation
In Table 4, we record the algorithm inference runtime for all the methods to evaluate their efficiency. Through analyzing the inference runtime of different methods, we have the following key observations:
(1) We can observe that our MAAdvisor’s algorithm runtime outperforms the prompt-based baseline LLMIdxAdvis in most of the settings. The reason is that with the multi-agent pipeline, our method’s only time overheads come from invoking the LLM API and refreshing column candidates, whereas LLMIdxAdvis faces an efficiency bottleneck due to the added cost of identifying the best demonstrations for each selection. In addition, LLMIdxAdvis suffers from high latency during its demonstration construction stage. It is reported that 17 hours and $90 are taken to collect the demonstrations (Zhao et al., 2025). It is also important to note that LLMIdxAdvis’ time cost, given increasing storage budgets, does not change a lot since they call the same demonstration match algorithm. In contrast, our MAAdvisor’s algorithm runtime may increase as the budget increases. It is because our index recommendation is storage-aware. The more budget given, the more local agents are planned by the planning agent. However, MAAdvisor’s time cost under the largest 50% budget is still lower than LLMIdxAdvis.
(2) The learning-based methods have significantly faster inference time compared to the other methods. The reason is that the learning-based methods only need to call the trained models with low latency during inference. However, in a zero-shot situation, the training time of the learned models shall also be considered as part of the algorithm latency. As an example, SWIRL’s training times for the TPC-H, TPC-DS, and JOB datasets are 1226.136, 3373.32, and 3635.00 seconds, respectively. In addition, both learned methods require an initial training set of data collected from the database, which is difficult to acquire in a zero-shot.
(3) Heuristic-based advisors on larger workloads generally suffer from much higher latency. The algorithm runtime for Relaxation on TPC-DS with an 83-query workload is more than 100 times the runtime on the TPC-H benchmark with 19 queries. For a larger workload and more complicated database, MAAdvisor presents a significant efficiency advantage. Although the DB2Advis baseline demonstrates a shorter runtime compared to our MAAdvisor, it trades off with a lower workload performance by using a simpler greedy optimization.
pipeline_example
(4) It is important to note that for the methods AutoAdmin and Drop following the DROP heuristic, the hyperparameter constraint for these methods is designed to be the maximum number of indexes to recommend. While under a storage budget constraint, the number of indexes to recommend is unknown; DBAs usually begin with a larger index number constraint and then locate the best one under the storage constraint. Therefore, the algorithm time we record, which is only the time for recommending the right index, is much shorter than the actual time required to obtain the final index recommendation from scratch. As an example, the total time of recommending indexes for TPC-DS under a 10% budget is 689.05 seconds for AutoAdmin and 5334.80 seconds for Drop.
We further evaluate the efficiency of our MAAdvisor with the LLM baseline LLMIdxAdvis in terms of monetary cost. The average number of input tokens and corresponding API cost for MAAdvisor and LLMIdxAdvis is also recorded in Table 5. It is observed that LLMIdxAdvis’ input lengths are more stable than MAAdvisor across benchmarks. The main reason is that LLMIdxAdvis uses a similar number of demonstrations in all benchmarks. In contrast, MAAdvisor’s input length depends on database complication and the number of column candidates. As an example, MAAdvisor takes 4 times more tokens for TPC-DS with 87 column candidates than TPC-H with 15 column candidates. Considering the complexity of the index recommendation task, MAAdvisor’s additional cost on more complicated workloads is totally acceptable.
7.4. Qualitative Evaluation
An example of MAAdvisor’s planning steps and index recommendation on TPC-H 20% budget can be found in Figure 7. We can observe that MAAdvisor efficiently recommends the indexes within five sub-steps, including selection and combination. Since the index selected is just within the storage budget, the revision agent is not called by the planning agent. It is also important to note that MAAdvisor adopts the same storage cost estimation as the other baselines to increase efficiency. Therefore, the actual storage cost of the output is smaller than the estimation in the planning phase. MAAdvisor can also opt to construct and calculate the actual index storage cost at each sub-step at a cost of a higher algorithm runtime.
In order to further analyze how well MAAdvisor addresses the index regression problem, we pick out the queries in each benchmark having any methods’ recommended indexes under 20% budget leading to a more than one-second’s regression. As shown in Figure 6, it is observed that DB2Advis has the most regression cases, trading off the high algorithm efficiency. Most heuristic-based advisors suffer from regressions on the most complicated DSB benchmark, since their optimization relies solely on the optimizer’s estimations. It is important to note that, at the workload level, significant improvements on some queries with a cost of regression on some other queries is acceptable as long as the overall execution time is decreased. On the other hand, although MAAdvisor does not result in the highest query time improvement on all mentioned queries, the revision agent manages to avoid significant regressions and therefore leads to a higher workload-level performance.
8. Ablation Studies
In this section, we first conduct an ablation study to evaluate the effectiveness of MAAdvisor’s workload representation technique and multi-agent pipeline. We will then analyze MAAdvisor’s generalisation ability over the LLM backbone and workload scale.
8.1. Effectiveness of Workload Representation
To validate the effectiveness of MAAdvisor’s workload representation step, we perform an ablation study in which the database schema and raw workload queries are directly input to the LLM agents. Instead of selecting from column candidates, the selection agent is now prompted to select one possible column index from the SQL queries. The result for this raw workload setting is presented in Table 6. We can observe that the performance drops significantly from 61.33% to 29.52%, which is reasonable since the planning and selection agents will plan and select indexes solely based on their domain knowledge and even without any demonstrations. In addition, the input context for the agents will be much longer, and LLMs’ hallucination and misunderstanding will increase significantly.
| Avg. Relative workload improvement | (%) |
| MAAdvisor (TPC-H 1x) | 61.33 |
| - Raw Workload | 29.52 |
| - One-call Recommendation | 11.20 |
| - Raw Workload and One-call Recommendation | 10.58 |
| - w/o Regression Indicator | 58.21 |
| MAAdvisor (GPT-3.5-turbo) | 60.01 |
| MAAdvisor (QWen3 32b) | 60.34 |
| MAAdvisor (QWen3 14b) | 40.86 |
| MAAdvisor (DeepCoder 14b) | 40.27 |
| MAAdvisor (TPC-H 20x) | 61.31 |
8.2. Effectiveness of Multi-agent pipeline
To explore the effectiveness of our multi-agent pipeline, we prompt the LLM to directly select from the column candidates and construct indexes in one step. Given the column candidates and storage budget, LLM is called only once to output all the recommended indexes. We present the one-call recommendation result in Table 6.
We can observe that the one-call recommendation yields an even lower performance of 11.20%. LLM fails to select effective indexes and control the storage budget at the same time. LLM tends to select fewer indexes with little improvement, but also little storage cost. In addition, we design another ablation with both one-call recommendation and raw workload input. The relative workload improvement further drops to 10.58%, showing that both our workload representation technique and the multi-agent pipeline are important in solving the index recommendation problem.
8.3. Effectiveness of Regression Indicator
We further test the effectiveness of the specially trained regression indicator embedded in the revision agent. We apply the MAAdvisor to the TPC-H benchmark with the regression indicator removed. The revision agent will then only rely on the provided expert experience to revise the current index recommendation. We can observe that MAAdvisor’s performance only suffers from a small drop. The main reason is that for smaller budgets, the index recommendation with a multi-agent pipeline is straightforward without too many revisions required, as shown in Figure 7. In addition, the revision agent alone can handle most of the severely regressive indexes by analysing the index utility estimation and learning from the expert experiences.
8.4. Different LLM Backbones
To evaluate MAAdvisor’s performance given different LLM backbones, we adopt one LLM API GPT-3.5-turbo (Brown et al., 2020), and three open-source LLMs, namely Qwen3-32B (Yang et al., 2025), Qwen3-14B, and DeepCoder-14B (Luo et al., 2025). DeepCoder-14B is the latest LLM that is specially finetuned on code-related data and achieves SOTA performance in code-related tasks. The results of adopting different LLM backbones are shown in Table 6. Since MAAdvisor relies on the reasoning ability of the LLM, we can observe that GPT-3.5-turbo and Qwen3-32B have comparable performance to GPT-4o-0806 and outperform Qwen3-14B. Smaller LLMs can also perform index recommendation following our MAAdvisor pipeline, but with a lower performance. It is important to note that the DeepCoder, specially fine-tuned on code-related data, underperforms the Qwen3 model with the same parameter size, showing that the coding ability of LLMs is not a must, given the rich information provided by the column candidates. In contrast, the reasoning ability of the LLMs is more important.
8.5. Different Workload Scales
In real-world applications, the workloads may always be updating, and such a situation is usually referred to as a dynamic setting. In order to simulate the change of scales, we construct another TPC-H workload of 380 queries named TPC-H 20x, where 20 random queries are generated for each query template. As shown in Table 6, MAAdvisor’s performance is robust against different workload scales. In addition, since the number of column candidates for the TPC-H 20x is almost the same as TPC-H 1x, the algorithm runtime for TPC-H 20x only increases from an average of 88.80 seconds to 92.91 seconds.
9. Conclusion and Discussion
In this work, we present MAAdvisor, a novel zero-shot index advisor that leverages the reasoning power of large language models through a hierarchical multi-agent pipeline. By representing the workload using column candidates and decomposing the index recommendation task into coordinated sub-steps handled by specialized agents, MAAdvisor achieves superior performance, strong generalisability, and high efficiency without the need for database-specific training. Extensive experiments across diverse benchmarks demonstrate that MAAdvisor not only outperforms traditional heuristic and learning-based approaches but also surpasses existing prompt-based LLM methods.
It is also important to acknowledge the current limitations and outline potential directions for future work. The efficiency of MAAdvisor may be further improved through the design of more precise and effective instruction sets. Additionally, fine-tuning the agent reasoning pipeline (Wang et al., 2024a) offers a promising avenue to reduce planning overhead and facilitate deployment on local LLMs. Furthermore, by capitalizing on the advanced reasoning and generalization capabilities of LLMs, the proposed multi-agent framework could be extended to address the index recommendation problem across heterogeneous DBMS environments and diverse SQL dialects.
Acknowledgements.
This work was supported by DAMO Academy through the DAMO Academy Research Intern Program.References
- (1)
- dsb ([n.d.]) [n.d.]. DSB Benchmark. https://github.com/microsoft/dsb/tree/main.
- hyp ([n.d.]) [n.d.]. HypoPG Extension. https://github.com/HypoPG/hypopg.
- Pos ([n.d.]) [n.d.]. PostgreSQL. https://www.postgresql.org.
- TPC ([n.d.]) [n.d.]. TPC-H Toolkit. https://www.tpc.org/tpc_documents_current_versions/current_specifications5.asp.
- Achiam and et al. (2023) Josh Achiam and Steven Adler et al. 2023. GPT-4 Technical Report. CoRR abs/2303.08774 (March 2023). https://arxiv.org/abs/2303.08774
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. CoRR abs/2005.14165 (May 2020). https://arxiv.org/abs/2005.14165
- Bruno and Chaudhuri (2005) Nicolas Bruno and Surajit Chaudhuri. 2005. Automatic physical database tuning: a relaxation-based approach. In Proceedings of the 2005 ACM SIGMOD International Conference on Management of Data (Baltimore, Maryland) (SIGMOD ’05). Association for Computing Machinery, New York, NY, USA, 227–238. https://doi.org/10.1145/1066157.1066184
- Chaudhuri and Narasayya (2020) Surajit Chaudhuri and Vivek Narasayya. 2020. Anytime Algorithm of Database Tuning Advisor for Microsoft SQL Server. (June 2020). https://www.microsoft.com/en-us/research/publication/anytime-algorithm-of-database-tuning-advisor-for-microsoft-sql-server/
- Chaudhuri and Narasayya (1997) Surajit Chaudhuri and Vivek R. Narasayya. 1997. An Efficient Cost-Driven Index Selection Tool for Microsoft SQL Server. In Proceedings of the 23rd International Conference on Very Large Data Bases (VLDB ’97). Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 146–155.
- Choenni et al. (1993) S. Choenni, H. Blanken, and T. Chang. 1993. Index selection in relational databases. In Proceedings of ICCI’93: 5th International Conference on Computing and Information. 491–496. https://doi.org/10.1109/ICCI.1993.315323
- Ding et al. (2021) Bailu Ding, Surajit Chaudhuri, Johannes Gehrke, and Vivek Narasayya. 2021. DSB: A Decision Support Benchmark for Workload-Driven and Traditional Database Systems. In VLDB 2022. https://www.microsoft.com/en-us/research/publication/dsb-a-decision-support-benchmark-for-workload-driven-and-traditional-database-systems/
- Ding et al. (2019) Bailu Ding, Sudipto Das, Ryan Marcus, Wentao Wu, Surajit Chaudhuri, and Vivek R. Narasayya. 2019. AI Meets AI: Leveraging Query Executions to Improve Index Recommendations. In Proceedings of the 2019 International Conference on Management of Data (Amsterdam, Netherlands) (SIGMOD ’19). Association for Computing Machinery, New York, NY, USA, 1241–1258. https://doi.org/10.1145/3299869.3324957
- He et al. (2016) Ji He, Jianshu Chen, Xiaodong He, Jianfeng Gao, Lihong Li, Li Deng, and Mari Ostendorf. 2016. Deep Reinforcement Learning with a Natural Language Action Space. In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Katrin Erk and Noah A. Smith (Eds.). Association for Computational Linguistics, Berlin, Germany, 1621–1630. https://doi.org/10.18653/v1/P16-1153
- Jha et al. (2025) Kunal Jha, Wilka Carvalho, Yancheng Liang, Simon Shaolei Du, Max Kleiman-Weiner, and Natasha Jaques. 2025. Cross-environment Cooperation Enables Zero-shot Multi-agent Coordination. In Forty-second International Conference on Machine Learning. https://openreview.net/forum?id=zBBYsVGKuB
- Ji et al. (2023) Ziwei Ji, Nayeon Lee, Rita Frieske, Tiezheng Yu, Dan Su, Yan Xu, Etsuko Ishii, Ye Jin Bang, Andrea Madotto, and Pascale Fung. 2023. Survey of Hallucination in Natural Language Generation. Comput. Surveys 55, 12 (March 2023), 1–38. https://doi.org/10.1145/3571730
- Kossmann et al. (2022) Jan Kossmann, Alexander Kastius, and Rainer Schlosser. 2022. SWIRL: Selection of Workload-aware Indexes using Reinforcement Learning. In Proceedings of the 25th International Conference on Extending Database Technology (EDBT). 2:155– 2:168.
- Lan et al. (2020) Hai Lan, Zhifeng Bao, and Yuwei Peng. 2020. An Index Advisor Using Deep Reinforcement Learning. In Proceedings of the 29th ACM International Conference on Information & Knowledge Management (Virtual Event, Ireland) (CIKM ’20). Association for Computing Machinery, New York, NY, USA, 2105–2108. https://doi.org/10.1145/3340531.3412106
- Leis et al. (2015) Viktor Leis, Andrey Gubichev, Atanas Mirchev, Peter A. Boncz, Alfons Kemper, and Thomas Neumann. 2015. How Good Are Query Optimizers, Really? Proc. VLDB Endow. 9 (2015), 204–215. https://api.semanticscholar.org/CorpusID:7953847
- Li et al. (2025) Ao Li, Yuexiang Xie, Songze Li, Fugee Tsung, Bolin Ding, and Yaliang Li. 2025. Agent-Oriented Planning in Multi-Agent Systems. In The Thirteenth International Conference on Learning Representations. https://openreview.net/forum?id=EqcLAU6gyU
- Liu et al. (2025) Yuchi Liu, Jaskirat Singh, Gaowen Liu, Ali Payani, and Liang Zheng. 2025. Towards Hierarchical Multi-Agent Workflows for Zero-Shot Prompt Optimization. In Workshop on Reasoning and Planning for Large Language Models. https://openreview.net/forum?id=RVvXOrP2qm
- Luo et al. (2025) Michael Luo, Sijun Tan, Roy Huang, Ameen Patel, Alpay Ariyak, Qingyang Wu, Xiaoxiang Shi, Rachel Xin, Colin Cai, Maurice Weber, Ce Zhang, Li Erran Li, Raluca Ada Popa, and Ion Stoica. 2025. DeepCoder: A Fully Open-Source 14B Coder at O3-mini Level. https://pretty-radio-b75.notion.site/DeepCoder-A-Fully-Open-Source-14B-Coder-at-O3-mini-Level-1cf81902c14680b3bee5eb349a512a51. Notion Blog.
- Maas et al. (2011) Andrew L. Maas, Raymond E. Daly, Peter T. Pham, Dan Huang, Andrew Y. Ng, and Christopher Potts. 2011. Learning Word Vectors for Sentiment Analysis. In Proceedings of the 49th Annual Meeting of the Association for Computational Linguistics: Human Language Technologies. Association for Computational Linguistics, Portland, Oregon, USA, 142–150. http://www.aclweb.org/anthology/P11-1015
- Paludo Licks et al. (2020) Gabriel Paludo Licks, Julia Colleoni Couto, Priscilla de Fátima Miehe, Renata de Paris, Duncan Dubugras Ruiz, and Felipe Meneguzzi. 2020. SmartIX: A database indexing agent based on reinforcement learning. Applied Intelligence 50, 8 (Aug. 2020), 2575–2588. https://doi.org/10.1007/s10489-020-01674-8
- Perera et al. (2021) R. Malinga Perera, Bastian Oetomo, Benjamin I. P. Rubinstein, and Renata Borovica-Gajic. 2021. DBA bandits: Self-driving index tuning under ad-hoc, analytical workloads with safety guarantees. In 37th IEEE International Conference on Data Engineering, ICDE 2021, Chania, Greece, April 19-22, 2021. IEEE, 600–611. https://doi.org/10.1109/ICDE51399.2021.00058
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics. https://arxiv.org/abs/1908.10084
- Sadri et al. (2020a) Zahra Sadri, Le Gruenwald, and Eleazar Lead. 2020a. DRLindex: deep reinforcement learning index advisor for a cluster database. In Proceedings of the 24th Symposium on International Database Engineering & Applications (Seoul, Republic of Korea) (IDEAS ’20). Association for Computing Machinery, New York, NY, USA, Article 11, 8 pages. https://doi.org/10.1145/3410566.3410603
- Sadri et al. (2020b) Zahra Sadri, Le Gruenwald, and Eleazar Leal. 2020b. Online Index Selection Using Deep Reinforcement Learning for a Cluster Database. In 2020 IEEE 36th International Conference on Data Engineering Workshops (ICDEW). 158–161. https://doi.org/10.1109/ICDEW49219.2020.00035
- Schlosser et al. (2019) Rainer Schlosser, Jan Kossmann, and Martin Boissier. 2019. Efficient Scalable Multi-attribute Index Selection Using Recursive Strategies. In 2019 IEEE 35th International Conference on Data Engineering (ICDE). 1238–1249. https://doi.org/10.1109/ICDE.2019.00113
- Schulman et al. (2017) John Schulman, Filip Wolski, Prafulla Dhariwal, Alec Radford, and Oleg Klimov. 2017. Proximal Policy Optimization Algorithms. CoRR abs/1707.06347 (July 2017). https://arxiv.org/abs/1707.06347
- Shen et al. (2023) Yongliang Shen, Kaitao Song, Xu Tan, Dongsheng Li, Weiming Lu, and Yueting Zhuang. 2023. HuggingGPT: Solving AI Tasks with ChatGPT and its Friends in Hugging Face. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 38154–38180. https://proceedings.neurips.cc/paper_files/paper/2023/file/77c33e6a367922d003ff102ffb92b658-Paper-Conference.pdf
- Shi et al. (2022) Jiachen Shi, Gao Cong, and Xiao-Li Li. 2022. Learned Index Benefits: Machine Learning Based Index Performance Estimation. Proc. VLDB Endow. 15, 13 (Sept. 2022), 3950–3962. https://doi.org/10.14778/3565838.3565848
- Siddiqui et al. (2022) Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri. 2022. DISTILL: low-overhead data-driven techniques for filtering and costing indexes for scalable index tuning. Proc. VLDB Endow. 15, 10 (June 2022), 2019–2031. https://doi.org/10.14778/3547305.3547309
- Skelley (2000) Alan Skelley. 2000. DB2 Advisor: An Optimizer Smart Enough to Recommend its own Indexes. In Proceedings of the 16th International Conference on Data Engineering (ICDE ’00). IEEE Computer Society, USA, 101.
- Wang et al. (2024a) Junjie Wang, Mingyang Chen, Binbin Hu, Dan Yang, Ziqi Liu, Yue Shen, Peng Wei, Zhiqiang Zhang, Jinjie Gu, Jun Zhou, Jeff Z. Pan, Wen Zhang, and Huajun Chen. 2024a. Learning to Plan for Retrieval-Augmented Large Language Models from Knowledge Graphs. In EMNLP (Findings). 7813–7835. https://aclanthology.org/2024.findings-emnlp.459
- Wang et al. (2024b) Zijia Wang, Haoran Liu, Chen Lin, Zhifeng Bao, Guoliang Li, and Tianqing Wang. 2024b. Leveraging Dynamic and Heterogeneous Workload Knowledge to Boost the Performance of Index Advisors. Proc. VLDB Endow. 17, 7 (March 2024), 1642–1654. https://doi.org/10.14778/3654621.3654631
- Wei et al. (2023) Jerry Wei, Jason Wei, Yi Tay, Dustin Tran, Albert Webson, Yifeng Lu, Xinyun Chen, Hanxiao Liu, Da Huang, Denny Zhou, and Tengyu Ma. 2023. Larger language models do in-context learning differently. CoRR abs/2303.03846 (March 2023). https://arxiv.org/abs/2303.03846
- Yang et al. (2025) An Yang, Anfeng Li, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chang Gao, Chengen Huang, Chenxu Lv, Chujie Zheng, Dayiheng Liu, Fan Zhou, Fei Huang, Feng Hu, Hao Ge, Haoran Wei, Huan Lin, Jialong Tang, Jian Yang, Jianhong Tu, Jianwei Zhang, Jianxin Yang, Jiaxi Yang, Jing Zhou, Jingren Zhou, Junyang Lin, Kai Dang, Keqin Bao, Kexin Yang, Le Yu, Lianghao Deng, Mei Li, Mingfeng Xue, Mingze Li, Pei Zhang, Peng Wang, Qin Zhu, Rui Men, Ruize Gao, Shixuan Liu, Shuang Luo, Tianhao Li, Tianyi Tang, Wenbiao Yin, Xingzhang Ren, Xinyu Wang, Xinyu Zhang, Xuancheng Ren, Yang Fan, Yang Su, Yichang Zhang, Yinger Zhang, Yu Wan, Yuqiong Liu, Zekun Wang, Zeyu Cui, Zhenru Zhang, Zhipeng Zhou, and Zihan Qiu. 2025. Qwen3 Technical Report. CoRR abs/2505.09388 (May 2025). https://arxiv.org/abs/2505.09388
- Zhang et al. (2023) Yue Zhang, Yafu Li, Leyang Cui, Deng Cai, Lemao Liu, Tingchen Fu, Xinting Huang, Enbo Zhao, Yu Zhang, Yulong Chen, Longyue Wang, Anh Tuan Luu, Wei Bi, Freda Shi, and Shuming Shi. 2023. Siren’s Song in the AI Ocean: A Survey on Hallucination in Large Language Models. CoRR abs/2309.01219 (Sept. 2023). https://arxiv.org/abs/2309.01219
- Zhao et al. (2025) Xinxin Zhao, Haoyang Li, Jing Zhang, Xinmei Huang, Tieying Zhang, Jianjun Chen, Rui Shi, Cuiping Li, and Hong Chen. 2025. LLMIdxAdvis: Resource-Efficient Index Advisor Utilizing Large Language Model. CoRR abs/2503.07884 (March 2025). https://doi.org/10.48550/arXiv.2503.07884
- Zhao et al. (2022) Yue Zhao, Gao Cong, Jiachen Shi, and Chunyan Miao. 2022. QueryFormer: a tree transformer model for query plan representation. Proc. VLDB Endow. 15, 8 (April 2022), 1658–1670. https://doi.org/10.14778/3529337.3529349
- Zhao et al. (2023) Yue Zhao, Zhaodonghui Li, and Gao Cong. 2023. A Comparative Study and Component Analysis of Query Plan Representation Techniques in ML4DB Studies. Proc. VLDB Endow. 17, 4 (Dec. 2023), 823–835. https://doi.org/10.14778/3636218.3636235
- Zhou et al. (2024) Wei Zhou, Chen Lin, Xuanhe Zhou, and Guoliang Li. 2024. Breaking It Down: An In-Depth Study of Index Advisors. Proc. VLDB Endow. 17, 10 (June 2024), 2405–2418. https://doi.org/10.14778/3675034.3675035
- Zhou et al. (2022) Xuanhe Zhou, Luyang Liu, Wenbo Li, Lianyuan Jin, Shifu Li, Tianqing Wang, and Jianhua Feng. 2022. AutoIndex: An Incremental Index Management System for Dynamic Workloads. In ICDE.
- Zhou et al. (2023) Yongchao Zhou, Andrei Ioan Muresanu, Ziwen Han, Keiran Paster, Silviu Pitis, Harris Chan, and Jimmy Ba. 2023. Large Language Models are Human-Level Prompt Engineers. In The Eleventh International Conference on Learning Representations. https://openreview.net/forum?id=92gvk82DE-

### v2

@@
-MAAdvisor: Zero-Shot Index Advisor using Multi-Agent LLMs
+AMAZe: A Multi-Agent Zero-shot Index Advisor for Relational Databases
 Abstract.
@@
 However, such methods still can not achieve the state-of-the-art (SOTA) results, and preparing the index selection demonstrations is also resource-intensive.
-To address these issues, we propose MAAdvisor, a zero-shot LLM-based index advisor with a multi-agent framework.
+To address these issues, we propose AMAZe, a zero-shot LLM-based index advisor with a multi-agent framework.
 We decompose the index recommendation problem into sub-steps, including planning, selection, combination, revision, and reflection.
 A set of LLM-embedded agents is designed to handle each one of the different sub-steps.
-Our method utilizes global agents to control the index selection process and local agents to select and revise indexes.
-Through extensive experiments, we show that our proposed MAAdvisor not only achieves the SOTA performance compared to the heuristic methods, but also outperforms learning-based and prompt-based methods with higher efficiency and better zero-shot inference ability.
+Our method utilizes high-level agents to control the index selection process and low-level agents to select and revise indexes.
+Through extensive experiments, we show that our proposed AMAZe not only achieves the SOTA performance compared to the heuristic methods, but also outperforms learning-based and prompt-based methods with higher efficiency and better zero-shot inference ability.
 1. Introduction
-Indexes in database systems significantly accelerate data retrieval and reduce redundant I/O operations by providing alternative access paths to data based on attribute values.
-With the appropriate index configuration, database systems can better support analytical workloads that are highly demanded by many modern applications, such as inventory management, online shopping, and real-time advertising.
+Indexes in database systems significantly accelerate data retrieval and reduce redundant I/O operations through providing alternative access paths to data based on attribute values.
+With the appropriate index configuration, database systems can better support analytical workloads that are highly demanded by many modern applications such as inventory management, online shopping and real-time advertisement.
 However, selecting and creating indexes always comes with significant computation time and storage cost.
 Therefore, given the constraints such as limited resource and budget, it is important for database administrators (DBAs) to select the optimal indexes that can meet the data management needs.
-| Heuristic-based method | Learning-based method | Prompt-based method | MAAdvisor (Ours) | |
+| Heuristic-based method | Learning-based method | Prompt-based method | AMAZe (Ours) | |
 | Zero-shot Generalization | Yes | No | No | Yes |
@@
 | Efficiency | Low | Low | High | High |
-To assist DBAs with index selection, index advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Lan et al., 2020; Perera et al., 2021; Sadri et al., 2020a, b; Skelley, 2000; Wang et al., 2024b; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997; Kossmann et al., 2022; Zhao et al., 2025; Zhou et al., 2022; Paludo Licks et al., 2020) are often adopted to automatically select the optimal indexes for a given query workload on the target database.
-Existing index advisors fall into three main categories: heuristic-based advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Skelley, 2000; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997), learning-based advisors (Lan et al., 2020; Perera et al., 2021; Sadri et al., 2020a, b; Wang et al., 2024b; Kossmann et al., 2022; Zhou et al., 2022; Paludo Licks et al., 2020), and prompt-based LLM advisors (Zhao et al., 2025).
+To assist DBAs with index selection, index advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Lan et al., 2020; Perera et al., 2021a; Sadri et al., 2020a, b; Skelley, 2000; Wang et al., 2024b; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997; Kossmann et al., 2022; Zhao et al., 2025; Zhou et al., 2022; Paludo Licks et al., 2020; Perera et al., 2021b; Wu et al., 2022; Wang and Yoneki, 2024; Chang et al., 2024) are often adopted to automatically select the optimal indexes for a given query workload in the target database.
+Existing index advisors fall into three main categories: heuristic-based advisors (Chaudhuri and Narasayya, 2020; Bruno and Chaudhuri, 2005; Skelley, 2000; Schlosser et al., 2019; Choenni et al., 1993; Chaudhuri and Narasayya, 1997), learning-based advisors (Lan et al., 2020; Perera et al., 2021a; Sadri et al., 2020a, b; Wang et al., 2024b; Kossmann et al., 2022; Zhou et al., 2022; Paludo Licks et al., 2020; Perera et al., 2021b; Wu et al., 2022; Wang and Yoneki, 2024; Chang et al., 2024), and prompt-based LLM advisors (Zhao et al., 2025).
 Heuristic-based advisors typically use rule-based or greedy algorithms to identify optimal indexes based on database optimizers’ cost estimations.
@@
 Moreover, the need for data annotations and model re-training poses a major limitation in data-sensitive domains like healthcare or security, where data collection and sharing are tightly regulated.
-To increase advisors’ generalisation and reduce recommendation time cost, prompt tuning on LLMs is adopted (Zhou et al., 2023).
-Such a method prompts LLMs with preprocessed index recommendation demonstrations and lets the LLM select columns to build the indexes.
-Although the generalisation and efficiency of advisors benefit from LLM, there are some limitations.
-First, the performance of the model is highly sensitive to the prompt design, such as the choice and format of demonstrations.
-Second, demonstrations on the analyzed database schema are required to achieve optimal results, therefore, preparing high-quality demonstrations is time-consuming and often requires expert domain knowledge.
-Lastly, inference with LLMs can be computationally expensive for large-scale workloads.
+To exploit the adaptability and semantic reasoning strengths of LLMs while mitigating both database optimizer dependency and data dependency, prompt tuning on LLMs (Zhou et al., 2023) is adopted.
+Such a method (Zhao et al., 2025) prompts LLMs with preprocessed index recommendation demonstrations and allows the LLM select columns to build the indexes.
+While this LLM-based approach enhances the generalization and efficiency of index advisors, it also presents notable limitations.
+Actually, prompt tuning approach exhibits limited effectiveness in zero-shot setting due to LLMs’ limited understanding of database-specific contexts and optimization nuances without concrete examples, necessitating few-shot learning with domain-specific demonstrations (Brown et al., 2020).
+However, preparing high-quality demonstrations is challenging and time-consuming.
+In addition, inference with LLMs can be computationally expensive for large-scale workloads.
+When processing complex workloads, the number of tokens to be encoded and reasoned over grows rapidly, leading to high memory consumption and long inference times.
+This computational overhead becomes particularly pronounced in online environments.
 Table 1 summarizes the characteristics and drawbacks of three types of advisors as discussed above.
-Given the limitations of existing methods, we aim to design an index advisor that addresses the following fundamental challenges in index recommendation: (1) Zero-shot Generalization: To accommodate the diverse landscape of real-world database deployments, an effective index advisor must exhibit strong generalization capabilities.
+Given the limitations of existing methods, we aim to design an index advisor that achieves the following key objectives in index recommendation: (1) Zero-shot Generalization: To accommodate the diverse landscape of real-world database deployments, an effective index advisor must exhibit strong generalization capabilities.
 Specifically, the ability to recommend optimal indexes for previously unseen databases and workloads without requiring retraining or fine-tuning on domain-specific data.
@@
 (3) Efficiency: The advisor must exhibit low computational overhead, enabling rapid index recommendations that scale to large workloads without imposing significant time costs.
-To address the challenges above, we propose our multi-agent index advisor based on LLM, named MAAdvisor.
-First, to address the zero-shot generalization challenge, we build upon the success of recent multi-agent work solving other problems under the zero-shot setting (Liu et al., 2025; Li et al., 2025; Jha et al., 2025).
-Multi-agent frameworks can enhance scalability by dividing complex problems into manageable components processed in parallel by specialized zero-shot agents.
-To address the difficulty of decomposing the index recommendation problem into sub-problems, we innovatively design a hierarchical multi-agent pipeline including global agents and local agents.
-Global agents decompose the problem into iterative sub-steps and plan the use of local agents in each sub-step.
-Local agents will then perform three key functions: (1) select candidate index, (2) revise the current index selection, and (3) construct composite indexes.
-This design exploits LLMs’ demonstrated reasoning capabilities in index recommendation and enables the system to generalize across diverse database schemas and workload patterns in zero-shot.
-Second, to solve the effectiveness challenge caused by the index regression problem in most advisors, we design two novel components for our index advisor: (1) a new workload representation paradigm for the local selection agent to accurately select optimal indexes, and (2) a dedicated local revision agent that identifies and eliminates potentially regressive index recommendations by analyzing estimation errors and performance trade-offs.
-Third, to guarantee the computational efficiency, we constrain the search space through a budget-aware planning mechanism where the global planning agent uses storage constraints as an optimization objective, effectively pruning the solution space explored by local agents.
+This efficiency is critical because database systems often handle massive datasets and complex query patterns where slow workload processing and index recommendation can delay essential optimizations, potentially impacting overall system performance and user experience.
+In order to achieve the objectives above, we propose AMAZe, a multi-agent zero-shot index advisor based on LLM.
+To equip AMAZe with zero-shot inference capability, we build upon recent advances in multi-agent frameworks that have demonstrated success in addressing other tasks under the zero-shot setting (Liu et al., 2025; Li et al., 2025; Jha et al., 2025).
+A multi-agent framework can better support zero-shot inference because it decomposes the complex index selection problem into specialized sub-tasks, allowing different agents to collaborate and compensate for the limitations of a single LLM.
+These sub-tasks enable the framework to leverage the general knowledge encoded in LLMs while structuring their reasoning in a way that aligns more closely with database-specific requirements.
+We further design three critical components within the multi-agent framework: efficient workload representation, budget-aware agent planning, and effectiveness-aware agent coordination.
+The efficient workload representation facilitates streamlined communication among agents and mitigates LLM hallucination by providing a concise yet informative summary of the workload.
+Complementarily, budget-aware agent planning enhances overall efficiency by guiding the LLM to select indexes that optimize resource usage while respecting storage and computation constraints.
+To tackle the index regression problem, which constitutes the primary barrier to achieving effectiveness, we design effectiveness-aware agent coordination.
+Through cooperative interactions, multiple agents iteratively refine the current index selection by evaluating index benefits, thereby ensuring a more robust, stable, and effective index recommendation pipeline.
 We conduct extensive experiments and recommend indexes for multiple popular datasets and workloads, including TPC-H, TPC-DS, DSB, and JOB.
-We observe that MAAdvisor’s benefit-to-cost ratio outperforms all the heuristic-based baselines, which are also zero-shot, in 10 out of the 16 settings.
-In addition, MAAdvisor significantly outperforms the learning-based methods, which require intensive training.
-Compared with the few-shot prompt-based baseline, MAAdvisor also achieves better performance with lower inference time and no domain-specific data.
+We observe that AMAZe’s benefit-to-cost ratio outperforms all the heuristic-based baselines, which are also zero-shot, in 10 out of the 16 settings, and matches the best of them in the remaining 6 settings.
+In addition, AMAZe significantly outperforms the learning-based methods which requires intensive training.
+Compared with the few-shots prompt-based baseline, AMAZe also achieves better performance with lower inference time without relying on domain-specific data.
 Our main contributions are as follows:
-To the best of our knowledge, we are the first to propose MAAdvisor, a zero-shot index advisor using LLMs, achieving SOTA results with higher efficiency and zero-shot generalization.
+We propose AMAZe, which achieves SOTA results with higher efficiency and zero-shot generalization. To the best of our knowledge, AMAZe is the first LLM-based zero-shot index advisor.
 We propose a new workload representation paradigm that is suitable for LLMs to understand and analyze the index recommendation problem.
 We design a multi-agent framework to solve the index recommendation problem.
-The problem is decomposed into multiple sub-steps, which are much easier to solve by a cooperation of the global and local agents.
+The problem is decomposed into multiple sub-steps, which are much easier to solve by a cooperation of the high-level and low-level agents.
+Extensive experiments on various workloads demonstrate that AMAZe achieves superior performance with lower inference time compared to baseline methods.
 2. Preliminary
-In this section, we will first clarify the concepts related to index recommendation in Section 2.1. Next, we will then explain the problem of an index advisor in Section 2.2.
-2.1. Workload, Index Candidate and Budget
-Workload: The workload is referred to as a set of queries that will be executed in the database.
+In this section, we will first clarify the concepts related to index recommendation in Section 2.1. Next, we will then formally define the problem of an index advisor in Section 2.2.
+2.1. Workload, Index Candidate and Constraints
+Workload: The workload is referred to as a set of queries that will be optimized in the database.
 The ultimate goal of the index recommendation problem is to optimize the execution efficiency of the entire workload instead of any specific query within it.
 Index Candidate: Most of the methods that solve the index recommendation problem start with collecting a set of index candidates to select from.
-Each candidate within this set usually stands for a table column queried by any of the queries in the workload.
-Storage Budget: Index selection in databases is subject to various practical constraints, most notably storage budget.
-Creating indexes consumes additional disk space and incurs extra cost during data modifications.
+Each candidate within this set usually stands for a table column accessed by any of the queries in the workload.
+Selection Constraints: Index selection in databases is subject to various practical constraints, such as index numbers and storage budget.
+Selection constraints in the paper is referred as the additional storage overhead (i.e., disk space) unless otherwise specified.
 During index selection, it is significant to balance performance gains against the increased resource usage, which necessitates prioritizing indexes that offer the highest benefit-to-cost ratio.
 2.2. Index Advisor
-Index advisors receive the workload of a specified database along with constraints like the maximum number of allowable indexes and storage budget, and then automatically generate an index recommendation.
+Index advisors receive the workload of a specified database along with selection constraints, and then automatically generate a set of recommended indexes.
 The functionality of an index advisor can be formally defined as follows:
 Definition 2.1.
-(Index Advisor): Given a database , a workload consisting of queries querying the database and a storage budget , an index advisor will automatically compute the column candidates and construct a set of indexes with columns , optimizing the cost of the workload by
-where calculates the execution time of the workload querying the database with the indexes created, and calculates the total storage cost to construct the indexes selected.
+(Index Advisor): Given a database , a workload consisting of a set of queries, and a storage budget , the index advisor problem seeks to automatically identify an optimal set of indexes that minimizes workload execution cost while respecting storage constraints.
+Formally, let denote the set of all possible column combinations that can be indexed in database .
+The index advisor solves the following optimization problem:
+where represents the power set of all possible column combinations, computes the total execution time of workload on database with index set , and calculates the total storage cost to construct the indexes selected.
 3. Related Work
-3.1. Index Advisors
-Heuristic-based Index Advisors: Heuristic-based advisors iteratively and greedily select candidate indexes until the budget constraint is met.
+3.1. Heuristic-based Index Advisors
+Heuristic-based advisors iteratively select candidate indexes for an input workload in a greedy fashion.
 Within each optimization step, the index advisors will update the index recommendation according to the estimated benefits of the index candidates.
-There are mainly two heuristics for the index advisors to follow: the ADD heuristic and the DROP heuristic.
-As an example of the ADD heuristic, Extend (Schlosser et al., 2019) starts with no indexes and adds one index at a time according to the highest benefit-to-cost ratio.
-In contrast, Drop (Choenni et al., 1993), following the DROP heuristic, starts with all index candidates and iteratively removes the least beneficial one until optimization constraints are met.
+There are mainly two heuristics for the index advisors to follow: ADD heuristic and DROP heuristic.
+As an example of ADD heuristic, Extend (Schlosser et al., 2019) starts with no indexes and add one index at a time according to the highest benefit-to-cost ratio until the budget constraint is met.
+In contrast, Drop (Choenni et al., 1993) following the DROP heuristic starts with all index candidates and iteratively removes the least beneficial one until only the desired number of indexes are left.
 In general, heuristic-based advisors generally take significantly longer computation time to recommend indexes, especially for databases with complicated schemas.
-In addition, heuristic-based methods rely solely on database optimizers to evaluate index benefits.
-According to (Ding et al., 2019), wrong estimations in column cardinality may mislead index advisors to select indexes that are not beneficial to the workload.
-Learning-based Index Advisors: While recent work (Zhou et al., 2024) shows that supervised learning approaches such as QueryFormer (Zhao et al., 2022) can be applied to index recommendation, state-of-the-art learning-based advisors (Kossmann et al., 2022; Wang et al., 2024b; Lan et al., 2020; Sadri et al., 2020a; Paludo Licks et al., 2020) predominantly adopt reinforcement learning (RL) algorithms.
+In addition, heuristic-based methods rely solely on database optimizers to evaluate index benefits, and wrong estimations in column cardinality may mislead index advisors to select indexes that are not beneficial to the workload (Ding et al., 2019).
+3.2. Learning-based Index Advisors
+State-of-the-art learning-based advisors (Kossmann et al., 2022; Wang et al., 2024b; Lan et al., 2020; Sadri et al., 2020a; Paludo Licks et al., 2020; Wu et al., 2022; Wang and Yoneki, 2024) predominantly adopt reinforcement learning (RL) algorithms for lower inference latency.
 For example, DQN (Lan et al., 2020) integrates Q-learning with deep neural networks to derive optimal policies from workload representations, while SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b) leverage the PPO (Schulman et al., 2017) algorithm for policy optimization, with BALANCE further employing a transfer learning framework to exploit historical workloads.
-RL-based advisors typically achieve much lower inference latency than heuristic-based counterparts; however, the training time for such methods cannot be overlooked.
-In addition, the generalisability of the RL-based methods is lower since training is required for each unseen database.
-Prompt-based Index Advisors: Large Language Models (LLMs) have recently emerged as a hot topic in machine learning research.
-These models have demonstrated a surprisingly strong ability to handle a variety of text-related tasks such as generation, decision-making, and reasoning.
-Therefore, (Zhao et al., 2025) proposes to utilize LLMs for index recommendation.
-They construct and offer in-context demonstrations for LLMs to learn from.
-LLMs will then directly recommend indexes for both met and unmet databases based on the demonstrations.
-When a new workload is given, LLMIdxAdvis explores the prepared demonstration pool and selects the optimal demonstration for the workload by cosine similarity.
-LLMIdxAdvis successfully addresses the problem of efficiency and robustness in index recommendation.
-However, the efficiency and generalisability are achieved at the cost of the effectiveness of the indexes recommended.
-LLMIdxAdvis achieves only comparable performance against other methods in simpler workloads and underperforms the SOTA methods given more complicated workloads and database schemas.
-3.2. Multi-agent Pipelines
-Leveraging the strong generalizing and reasoning abilities of LLMs, recent research proposes the idea of agents (Shen et al., 2023; He et al., 2016).
-LLM agents employ a more complex, modular architecture, enabling end-to-end problem-solving and the ability to handle multi-step tasks.
-In a standard multi-agent pipeline (Liu et al., 2025; Li et al., 2025; Jha et al., 2025), a complex problem is divided into different sub-problems.
-The sub-problems are assigned to specialized agents that collaborate in parallel.
-Each agent will focus on one task, such as data pre-processing, information retrieval, reasoning, or generation.
-Communication occurs through shared memory or message passing among different agents.
-The agents can be either rule-based or learning-based, which improves scalability and interpretability, enabling efficient problem-solving for complex workflows requiring coordination, specialization, and iterative refinement among agents.
-4. MAAdvisor Pipeline Overview
-Figure 1 illustrates the pipeline of MAAdvisor.
-In this section, we will explain our motivations in the method design and offer an overview of the two key sections in our method, named workload representation (4.1) and multi-agent pipeline (4.2).
+However, RL-based methods need significantly longer training time and have lower generalisability, as training is required for each unseen database.
+Alternative learned models are proposed to address the limitations.
+For example, DBA Bandits (Perera et al., 2021b) employs a Multi-Armed Bandits framework, where index selection policies are defined and refined through neural networks to enhance computational efficiency.
+AutoIndex (Zhou et al., 2022) leverages a Monte Carlo Tree Search algorithm, adopting an ADD-style strategy similar to heuristic-based methods for index selection.
+More recently, MFIX (Chang et al., 2024) introduces a compact tree-structured index space and improves search efficiency by utilizing a multi-fidelity Bayesian optimization (MFBO) algorithm.
+Despite these advances, such methods often suffer from poor generalization or trade-offs between efficiency and effectiveness.
+Moreover, they still require retraining when the underlying data changes and fail to operate in zero-shot scenarios.
+It is noteworthy that some recent work focuses on specific sub-tasks of index recommendation including workload compressing (Siddiqui et al., 2022a; Brucato et al., 2024), index candidate filtering (Siddiqui et al., 2022b) and index evaluation (Zhao et al., 2022).
+As these studies require considerable extra design to establish a complete index advisor, we leave them for future discussion.
+3.3. Prompt-based Index Advisors
+Recently, LLMIdxAdvis (Zhao et al., 2025) first utilizes LLMs to generate index suggestions.
+Specifically, it leverages a prepared demonstration pool and employs cosine similarity to select the most relevant in-context demonstrations for a given workload.
+While this strategy improves efficiency and generalizability in index selection, these gains come at the expense of recommendation quality.
+Specifically, LLMIdxAdvis struggles to handle complex workloads and database schemas due to its reliance on limited demonstrations.
+In addition, it progressively builds index in a single-agent, greedy manner without refinement or verification, leading to significant hallucination problems.
+Deviating from LLMIdxAdvis, our work explores a multi-agent framework, which doesn’t rely on demonstrations and mitigates hallucinations via cross-agent coordination.
+4. Framework Overview
+In this section, we will give an overview of how we design our AMAZe framework to address the current challenges for index recommendation problem.
+In Section 4.1, we will first explain how we solve the zero-shot generalization challenge using a multi-agent pipeline.
+In Section 4.2 we will then introduce our novel framework designs, namely the workload representation paradigm and effectiveness-aware agents for better effectiveness and efficiency.
 | Execution time (s) | With regressions | Remove regressions |
@@
 | GPT-4o | 186.70 | 155.54 |
-MAAdvisor
-4.1. Workload Representation
-Motivation: To enable zero-shot LLM reasoning effectively and efficiently, two main challenges lie within the information for index recommendation, namely long input context and lack of domain knowledge.
-First, the database information, like schema, workloads, and query plans, can be extremely long in real-world applications.
-For example, a workload consisting of 83 queries from the TPC-DS benchmark alone can take 36518 tokens’ length in GPT-4’s input.
-Not only do smaller LLMs may not have the ability to handle such long input sequences, but the cost and hallucination (Ji et al., 2023; Zhang et al., 2023) for larger LLMs dealing with long contexts can also be considerably large.
-Therefore, a new representation is required to summarize the database-related information into effective and precise LLM inputs.
-Second, although previous work (Wei et al., 2023) has proven LLMs’ strong generalization and reasoning ability given inputs similar or related to their training data, general-purpose LLMs always have insufficient training data related to data management and database indexing.
-Therefore, effective workload representation is vital for solving the index recommendation problem with LLMs.
-Unfortunately, the workload representation techniques from existing methods are not fit for LLMs.
-The learning-based methods usually encode the database and workload information into embedded features (Lan et al., 2020; Wang et al., 2024b).
-However, such features are too abstract for LLMs to understand and use as inputs.
-The prompt-based method (Zhao et al., 2025) also attempts to address this problem by extracting index-related information in the workload for the LLMs.
-However, the information is yet stored in densely structured nested dictionaries.
-It is difficult for LLMs to locate the information needed, and hallucination may be triggered according to our experiments in Section 7.2.
-Therefore, a workload representation technique that is able to both shorten the LLM input length and preserve the critical information is required for LLMs to perform the index recommendation task.
-Overview: To address the challenges above, we reformulate and summarize the inputs in a textual way that LLMs can easily understand and use.
-The information from workload, database schema, and optimiser that is relevant to index recommendation is extracted, merged, and grouped by the indexable column candidates.
-Given all the column candidates and their relevant information affecting index recommendation, the problem of using LLMs to construct indexes from raw workload can be simplified into selecting and combining columns from existing candidate choices.
-At the same time, by merging inputs on a per-column basis, the LLM’s prompt length is capped at the database’s total column count.
-In other words, regardless of how many queries appear in the workload, the set of candidate columns can never surpass the number of columns defined in the schema.
-4.2. Multi-agent Pipeline
-Motivation: To explore the capability of LLMs to solve the index recommendation task, we first conduct a pioneering experiment on the TPC-DS benchmark.
-We generate 100 small workloads, with each consisting of 3 to 10 random queries querying the TPC-DS database, and set a 2 GB storage budget.
-The results of directly letting GPT-4o (Achiam and et al., 2023) select indexes in a zero-shot manner given column candidates, and the heuristic-based baseline Extend (Schlosser et al., 2019) are presented in Table 2.
-It is observed that GPT-4o underperforms Extend by 25.42 seconds on average.
-Although introducing GPT-4o increases the scalability of the index recommendation, it demonstrates that the problem is yet too complicated for a single call of LLM to solve in zero-shot.
-Similar findings have been reported in (Zhao et al., 2025) that even with one-shot demonstration, LLMs still fail to outperform existing methods.
-However, when we remove the regressing indexes from the workloads and only keep the improving ones, it is observed that GPT-4o can outperform Extend.
-It demonstrates LLM’s potential in index recommendation and also justifies that an efficient and well-designed framework powered by LLM can successfully address the problem.
-Pioneer experiments also reveal that GPT-4o’s index recommendations fluctuate considerably, highlighting the need for a budget-aware planning pipeline to ensure the LLM produces consistent outputs while adhering to budget constraints.
-Overview: To address the challenges above, we propose to simplify the index recommendation problem using a hierarchical multi-agent pipeline with budget-aware planning.
-At the global level, the multi-agent pipeline reflectively decomposes the problem into LLM-solvable sub-steps, including selecting, revising, or combining column candidates.
-At the local level, each sub-step is handled by a specialized local agent embedded with LLM.
-According to language model theory (Brown et al., 2020), LLMs produce more accurate and consistent outputs when handling simpler tasks.
-By breaking down the complex problem into smaller, well-defined sub-tasks and delegating them to specialized LLM agents, our multi-agent pipeline achieves both stability and efficiency in index recommendation.
-Moreover, by estimating the storage consumption at each sub-step, subsequent steps are planned accordingly, enabling budget-aware index recommendation.
+AMAZe
+4.1. Multi-agent Based Pipeline
+Motivation.
+Although LLMs demonstrate strong potential for zero-shot generalization in index recommendation, it is hard to optimize the LLM prompts for each database schema.
+To investigate this limitation, we conduct a preliminary experiment on the TPC-DS benchmark.
+We construct 100 small workloads, each containing 3–10 randomly selected queries over the TPC-DS schema, under a fixed 2 GB storage budget.
+We compare the performance of GPT-4o (Achiam and et al., 2023), applied directly in a zero-shot manner with workload and query plan inputs, against the heuristic baseline Extend (Schlosser et al., 2019).
+Table 2 shows that GPT-4o underperforms Extend by an average of 15.8 % (25.42 seconds).
+While GPT-4o can generate reasonable index recommendations without additional training, its performance remains limited and inconsistent across complex workloads.
+These findings confirm that index recommendation is far too complex to be solved effectively with a single zero-shot LLM call, highlighting the need for a well-designed LLM-powered framework that can fully harness LLMs’ zero-shot generalization capabilities beyond direct prompting alone.
+Inspired by the recent advancements in multi-agent LLM pipelines that divide a complex problem into sub-problems handled by specialized agents (Liu et al., 2025; Li et al., 2025; Jha et al., 2025), we aim to decompose the index recommendation task into multiple sub-tasks like index selection and combination.
+This design enhances scalability, interpretability, and efficiency, particularly well-suited for complex workloads that demand coordination, specialization, and iterative refinement.
+Pipeline Overview.
+We introduce a solution based on a hierarchical multi-agent based index recommendation pipeline, illustrated in Figure 1.
+At the high-level, the pipeline reflectively decomposes the complex index recommendation task into a sequence of LLM-solvable sub-steps, such as selecting promising column candidates, revising suboptimal or redundant choices, and composing indexes when beneficial.
+At the low-level, each sub-step is handled by a specialized agent powered by an LLM, ensuring that decisions remain accurate and contextually appropriate.
+This design directly addresses the challenge that LLMs struggle with the full index recommendation problem in a single zero-shot call: the multi-agent pipeline excels at solving simpler, well-defined tasks with greater accuracy and consistency (Brown et al., 2020).
+By orchestrating multiple agents, our pipeline leverages the generalization power of LLMs while avoiding their weaknesses on overly complex reasoning.
+4.2. Key Component Designs
+Design Rationale.
+To develop an effective multi-agent framework, we must solve three critical design problems: (1) how to efficiently represent workloads for LLM agents, (2) how to incorporate storage budget constraints into the agent pipeline, and (3) how to prevent index regression that undermines recommendation quality.
+These design challenges are elaborated below:
+(1) Workload Representation.
+Jointly encoding database schemas, workloads, and query execution plans often exceeds the context lenght limitations of LLMs—the TPC-DS workload with 83 queries requires over 36,000 tokens, surpassing most LLM’s context capacity and incurring substantial costs.
+Beyond computational constraints, long contexts reduce agent scalability and decision accuracy.
+Moreover, existing workload representations are unsuitable for agent-based systems.
+Learning-based methods encode information into embedded features (Lan et al., 2020; Wang et al., 2024b) that agents cannot interpret, while prompt-based approaches (Zhao et al., 2025) encode each query into a nested dictionary and append them as a list input that agents struggle to navigate effectively.
+These verbose, code-format representations misalign with how LLM agents process and reason about information.
+Therefore, we need to design a specialized workload representation technique that enables efficient agent communication and reasoning.
+(2) Budget Constraints.
+Existing agent frameworks lack mechanisms to enforce storage constraints across multiple agents.
+Our preliminary experiments reveal that LLM does not always know when to stop the recommendation, demonstrating that agents operating without budget coordination produce inconsistent and resource-inefficient results.
+Thus, we must design budget-aware coordination mechanisms that enable agents to collaborate while respecting storage limitations.
+(3) Regression Prevention.
+The index regression problem—where recommended indexes degrade query performance-is amplified in multi-agent settings due to three factors: suboptimal decisions by individual agents, LLM hallucinations (Ji et al., 2023; Zhang et al., 2023), and compounding errors across agent interactions.
+As shown in Table 2, when regressing indexes are removed, GPT-4o surpasses traditional baselines, indicating that not only regression prevention is crucial, but also optimal index recommendation is achievable by LLMs.
+However, general-purpose LLMs lack database domain knowledge (Wei et al., 2023), leading agents to make uninformed decisions that cascade through the pipeline.
+Existing agent frameworks provide no mechanisms to detect, prevent, or correct such regression patterns.
+Consequently, we need to design agent coordination strategies that incorporate database-specific validation, decompose complex decisions into reliable subtasks, and implement feedback mechanisms to prevent regression propagation.
+Key designs. To address the aforementioned challenges, we propose three key designs that enable effective multi-agent coordination for index recommendation.
+(1) Efficient Workload Representation.
+We design a novel textual workload representation that transforms the complex index recommendation problem into an LLM-compatible format.
+As illustrated in Figure 1, our approach extracts only index-relevant information and reorganizes it at the column level, merging all workload details associated with each database column to construct a compact set of indexable candidates.
+This design provides two critical advantages.
+First, it ensures efficiency and scalability by bounding the input length to the number of database columns rather than the number of queries—making the representation size independent of workload complexity.
+Second, it reduces hallucination by reframing index recommendation as a structured selection task where agents choose and combine predefined indexable columns, aligning the problem with LLMs’ decision-making strengths rather than requiring free-form generation.
+(2) Budget-Aware Agent Planning.
+Our pipeline incorporates budget-aware coordination mechanisms that enable high-level agents to make resource-conscious decisions throughout the recommendation process.
+The planning agent receives the current storage consumption at its stage and communicates constraints to subsequent low-level agents, ensuring that the final recommendations remain feasible under real-world storage limitations.
+This index budget management prevents inefficient computations while maintaining the flexibility for agents to explore diverse optimization strategies within acceptable bounds.
+(3) Effectiveness-Aware Agent Coordination.
+To prevent index regression, we design specialized low-level agents targeting the two primary causes of performance degradation: LLM hallucination and database optimizer estimation errors.
+As shown in Figure 1, our coordination strategy employs two complementary agents: The combination agent addresses optimizer estimation bias by merging individually suboptimal indexes into composite indexes that better capture multi-column access patterns, effectively combining inefficient recommendations into more beneficial ones.
+The revision agent eliminates LLM-generated errors, including hallucinated indexes that do not exist in the schema, redundant duplicates, and indexes that empirically degrade performance.
+Importantly, the cooperation between these two agents not only addresses hallucination and misestimation but also simulates a hybrid of ADD and DROP index recommendation heuristics, where new composite indexes can be constructed while regressive ones are systematically removed.
+This coordinated workflow enables the system to leverage the strengths of both heuristics, leading to a more robust, stable, and effective index recommendation pipeline than relying on any single mechanism.
+agents
 5. Workload Representation
-In this section, we will introduce our workload representation technique.
+In this section, we will introduce in detail our efficient workload representation.
 In Section 5.1, we will first introduce the information we extract as a representation for each query in the workload.
-Then in Section 5.2, we will explain how we reformulate and merge the information in the whole workload into column candidates.
-5.1. Query Information for Index Selection
-To fully leverage the reasoning capabilities of LLMs, the information input should be concise and readily interpretable.
-However, the data extracted directly from SQL queries is often highly structured and domain-specific.
-Therefore, it is necessary to extract and reformulate the query-related information for the LLMs to interpret and process effectively.
-Building on insights from prior studies (Zhao et al., 2023, 2025; Siddiqui et al., 2022) and our preliminary experiments, we summarize key principles for extracting query-level information that are most relevant and effective for index recommendation with LLM:
-Principle 1: To perform zero-shot inference, information that varies with databases and does not significantly affect index selections, for example, the data values in join conditions, predicates, and database knobs, can be excluded from the input;
-Principle 2: Only columns that appear in any of the workload queries need to be considered as candidates to construct indexes, which will be called ‘indexable columns’.
-Based on the principles above, the information we extract from a query representing each indexable column is as follows:
-Table and column name: Table and column names are the key attributes for indexes.
-In addition, the semantics of the table and column names can usually offer the LLMs additional information during index recommendation.
-For simplification, the format ‘TableName.ColumnName’ will be used in the column candidate for LLMs to choose during the selection phase;
+Then in Section 5.2, we will explain how we further enhance pipeline efficiency by reformulating and merging the information in the whole workload into column candidates.
+5.1. Query Information Extraction
+To fully leverage the reasoning capabilities of LLMs, the input information should be concise and readily interpretable.
+However, data extracted directly from SQL queries is often highly structured and domain-specific, making it challenging for LLMs to process effectively.
+Therefore, it is necessary to extract and reformulate query-related information into a format that LLMs can interpret and process efficiently.
+Since we need to build index candidates derived from related columns, we first extract column information from workload queries.
+Building on insights from prior studies (Zhao et al., 2023, 2025; Siddiqui et al., 2022b) and our preliminary experiments, we identify that the essential column-related information includes:
+Indexable column names: Table and column names are the key attributes for constructing indexes.
+We define each column that appears in the workload as an ‘indexable column’.
+Specifically, the format ‘TableName.ColumnName’ will be used in the column candidate for LLMs to choose during the selection phase;
 Column operator type: Operators play important roles in index recommendation.
@@
 Generally, indexes on higher cardinality columns allow easier access to specific rows.
-However, indexes on low cardinality columns may also help in specific situations, for example, the column value queried is extremely infrequent compared to the other values in the same column;
+However, indexes on low cardinality columns may also be beneficial in specific situations, such as when the queried column value is extremely infrequent compared to others in the same column;
 Column distribution: Column distribution includes data histogram, minimum, and maximum values within the column.
@@
 Although the real storage cost can not be calculated before the actual index creation, the database optimizers can estimate the storage using hypothetical indexes (hyp, [n.d.]);
-Estimated column utility: Apart from the information above, an estimation of the potential improvement from a column index is vital for telling if a column candidate is worth selecting (Siddiqui et al., 2022).
+Estimated column utility: Apart from the information above, an estimation of the potential improvement from a column index is vital for telling if a column candidate is worth selecting (Siddiqui et al., 2022b).
 We define the utility of a potential index on a query as follows:
 Definition 5.1.
-(Index Utility): Given an index to create and a query , the utility of creating the index for is defined by query total cost without and with index per unit of storage cost:
-where the database optimizer uses to estimate the total time cost of the query , and uses to estimate the storage cost of the index .
+(Index Utility): Given an index and a query , the utility of creating the index for is defined by query total cost without and with index per unit of storage cost:
+where database optimisor uses to estimate the total time cost of the query , and uses to estimate the storage cost of the index .
 The utility of a column on query will then be calculated as .
 From the calculation, the estimated column utility is designed to be compound information, including column cardinality, query selectivity, and operations on the column.
-5.2. Column Candidate Generation
-In Section 5.1, we outlined the information required from each query for the index recommendation task.
-However, there still exist considerable duplications in information derived from each column.
-For a workload with indexable columns, where each column appears in queries on average, a total of pieces of information are collected.
-To enable the LLM to provide effective selections, this information must be consolidated so that each candidate item corresponds to a distinct column.
-It is observed that, among the information collected for a column, only the column operator type and the estimated column utility vary across queries.
-The column operator type can easily be merged by recording all the operators in a list.
-Consequently, an algorithm merging the estimated column utility for all queries in the workload can effectively eliminate duplicate column candidates.
-Since the goal of index recommendation is to improve workload performance, the index that reduces the execution cost more should be considered as more beneficial.
-Therefore, we merge the estimated column utility for the same column in different queries by aggregating their utility values.
-If a column has a higher utility, it means that selecting the column as an index will result in an estimated higher workload performance gain.
-For the same workload above, there will be exactly column candidates left for the indexable columns after the merging.
-With the utility merged by the columns, each column candidate is represented as a dictionary of information so that the LLMs can easily access and read.
-It is also important to note that the merging also benefits the LLM efficiency, especially for larger workloads, by largely reducing the input length.
-agents
+5.2. Workload Information Summarization
+Motivation.
+Given a workload with queries, we could extract column candidates by individually processing each query and extracting an average of indexable columns per query.
+However, this approach is inefficient because many columns appear repeatedly across different queries, creating numerous duplicates.
+To enable the LLM to make effective and efficient index selections, we must consolidate this information so that each candidate corresponds to a unique column rather than being repeated across multiple queries.
+To achieve this goal, we aggregate information for columns with the same name, ensuring that our final candidate set contains only unique columns.
+For most column attributes, consolidation is straightforward.
+The estimated column cardinality, column distribution, and estimated column storage remain constant for a given column regardless of which query references it, so we retain these values directly.
+However, two attributes vary across queries and require special handling:
+(1) column operator types: We merge these by collecting all operators observed for a column across different queries into a single comprehensive list.
+(2) estimated column utility: This requires a more principled aggregation approach.
+Since the goal of index recommendation is to enhance overall workload performance, columns that provide greater reductions in query execution cost should be prioritized.
+Therefore, we aggregate utility values across all queries that reference a given column.
+A higher aggregated utility indicates that indexing this column will likely yield greater overall performance improvement for the entire workload.
 6. Multi-agent Pipeline
-In this section, we will introduce our multi-agent pipeline in which the index recommendation problem is decomposed into sub-steps for agents to solve.
-As shown in Figure 2, the pipeline consists of two types of agents: local agents and global agents.
-In Section 6.1, we will first introduce the three local agents that will select, combine, or revise the index recommendation in each sub-step.
-We will then explain our design for the two global agents planning each sub-step and reflecting on the plans in Section 6.2.
-In Section 6.3, we will illustrate the whole multi-agent pipeline on how we let the agents cooperate and communicate to perform index recommendation.
-6.1. Local Agents
-Selection agent: As the key function of our index advisor, the selection agent plays the role of selecting new columns from the column candidates given.
-The inputs for the selection agent include previous selected indexes, column candidates to select from, total storage budget, and the storage budget already used.
-In each single call of the selection agent, LLM is prompted with the inputs to select one new column into the current index recommendation as a single-column index.
-Combination Agent: As noted earlier, the selection agent produces only single-column indexes.
-However, when multiple columns from the same table are frequently queried together, a composite index on those columns may deliver better performance than maintaining them as separate single-column indexes.
-To address this, we design a combination agent that can merge such indexes.
-To identify optimal composite indexes, we follow the approach in (Siddiqui et al., 2022), collecting potential rules for combining single-column indexes.
-As illustrated in Figure 3(b), these rules specify possible column combinations and their ordering based on query operations, which can be readily extracted from the queries themselves.
-In addition, we incorporate external knowledge from diverse sources such as the Internet, expert practices, and textbooks to fully leverage the LLM’s reasoning capabilities, thereby infusing domain expertise and facilitating effective index combination.
-Examples of external knowledge can also be found in Figure 3(c).
-Given the input of current selected indexes with operations, index combination rules, and external knowledge, the combination agent will decide if some indexes can be combined into a composite index and output the new index recommendation.
+This section presents our hierarchical multi-agent pipeline for index recommendation, decomposing the problem into structured sub-steps handled by specialized agents.
+Section 6.1 introduces three low-level agents that select, combine, and revise indexes within individual sub-steps.
+Section 6.2 details two high-level agents responsible for planning and resource allocation under budget constraints.
+Section 6.3 demonstrates how these agents interact to achieve budget-aware planning and coordination.
+6.1. Low-level Agents
+Selection agent: As the key function of index advisor, the selection agent plays the role of selecting new columns from the column candidates given.
+As shown in Figure 2, the inputs for the selection agent include previous selected indexes, column candidates to select from, total storage budget, and the storage budget already used.
+Once the selection agent is called, LLM is prompted to select one new column into the current index recommendation as a single-column index.
+Combination Agent: As noted earlier, the selection agent produces only single-column indexes, which are often insufficient when queries involve multiple columns from the same table.
+In such cases, composite indexes can provide greater efficiency than maintaining several separate single-column indexes.
+To address this, we introduce the combination agent, which merges selected single-column indexes into composite indexes that better capture multi-column access patterns.
+In doing so, it also mitigates optimizer estimation bias by transforming individually suboptimal indexes into beneficial composite ones.
+To guide this process, we first adopt the approach in (Siddiqui et al., 2022b), which outlines rules for combining indexes, including valid column groupings and orderings inferred from query structures, as illustrated in Figure 3(b).
+Beyond combination rules, we further enhance the agent with external knowledge from sources such as textbooks, expert practices, and the Internet, enabling it to incorporate domain expertise into index construction.
+Examples of such knowledge are shown in Figure 3(c).
+As shown in Figure 2, given the current indexes, query operations, combination rules, and external knowledge, the combination agent determines whether certain indexes should be merged, and outputs updated recommendations.
 prompt examples
-Revision Agent: The revision agent will be adopted to eliminate the potential duplicates incurred and index regressions caused by the current index recommendation.
-The duplicates can be easily removed by simple coding techniques.
-However, as addressed by the pioneer experiments in Table 2, index regression is one of the unsolved problems for both traditional and LLM-based index advisors.
-One of the reasons for the problem is the database optimizer’s estimation error for cardinality, which is used for evaluating the index’s potential benefits.
-Apart from the optimizer estimations, LLMs can also obtain cardinality-related information from the name and distribution of the columns.
+regression_indicator
+Revision Agent: The revision agent will be adopted to eliminate the hallucinations and potentially inefficient indexes in the current index recommendation, which may lead to index regression.
+First, a hallucination check will be conducted in the revision agent to identify and correct the wrong or non-existing index selections.
+Second, to mitigate the impact of potentially inaccurate estimations made by the database optimizer, a problem that is well documented even in traditional optimization pipelines, the LLM is guided by a three-fold justification mechanism that draws upon column information, expert experience and regression indications, as shown in Figure 2.
+Apart from the optimizer estimations, LLMs can also obtain cardinality-related column information from the name and distribution of the candidate columns selected.
 For example, the cardinality for the column ‘country_name’ is highly likely to be smaller than the column ‘order_id’.
 Therefore, we prompt the LLM to identify and remove columns from the selected index when a significant discrepancy exists between the estimated cardinality and the LLM’s analysis.
-To enhance the reasoning, we incorporate expert knowledge drawn from diverse sources such as common practices and online discussions.
+To enhance LLM’s ability to tell a regressing index, we inject expert knowledge drawn from diverse sources such as common practices and online discussions.
 Insights of inefficient indexes are then distilled into expert experience rules, which guide the detection of regressive indexes.
@@
 To strengthen the LLM’s capacity for independent and robust reasoning beyond reliance on the expert experiences, we also adopt a regression indicator to provide the LLM with an objective, data-driven measure that complements LLM reasoning.
-Regression indicator: Prior research (Ding et al., 2019; Shi et al., 2022) has attempted to address the index regression prediction problem using deep learning techniques.
-They label query plan pairs before and after index creation as either ‘improvement’ or ‘regression’, and train models on these labeled examples to forecast whether a given index will degrade query performance.
-However, these methods have limitations in handling zero-shot index recommendations at the workload level.
-Specifically, they focus only on query-level regressions rather than considering the impact on the entire workload.
-Moreover, their models require retraining for each new dataset, resulting in limited robustness and generalizability.
-To address the aforementioned limitations, we devise a workload-level regression indicator that is trained on features invariant to specific databases and workloads.
-The overall architecture is depicted in Figure 4.
+Regression indicator: Since prior learned index regression indicators (Ding et al., 2019; Shi et al., 2022) are dataset-specific and only detect query-level regressions, we propose a workload-level regression indicator which is trained using information invariant to databases and workloads.
+The structure of the model is presented in Figure 4.
 The workload’s query plans, both with and without the indexes, are extracted as inputs for the indicator.
@@
 The workload representation is constructed by concatenating the root vectors of all queries.
-Finally, given the pair of workload vectors before and after index creation, we employ an MLP classifier to predict potential index regressions from -1 to 1 based on the difference between the two representations.
-Each index will be assigned a predicted regressive score that the LLM can use as a reference.
+Finally, given the pair of workload vectors before and after index creation, we employ an MLP classifier to predict potential index regressions based on the difference between the two representations.
+Each index will be assigned a regressive score predicted that the LLM can use as a reference.
 Inspired by (Zhao et al., 2023), we extract only the essential features from each tree node that influence index performance, including the operator type, table, column, estimated cardinality, and cost values.
@@
 To enhance the robustness of the regression indicator, we employ Sentence Transformers (Reimers and Gurevych, 2019) to generate database-invariant embeddings for the table and column information.
-regression_indicator
-6.2. Global Agents
-Planning agent: The planning agent plays an important role in the multi-agent pipeline for managing problem sub-steps and determining which local agents to deploy at each step.
-Given the column candidates, current indexes, total storage budget, and budget already used, the planning agent first evaluates the stop criteria.
-If these criteria are not met, it outputs an action for the current sub-step: ’Selection’, ’Combination’, or ’Revision’, each invoking a corresponding local agent.
-When the planning agent determines that the index selection is optimized or that no additional columns can be chosen within the budget, it triggers the ’Stop’ action to conclude the optimization.
-Reflection agent: In practical use, it has been observed that planning agent’s strategies for local agents can be inefficient at times, as it may suffer from hallucinations or become trapped in local optima.
+6.2. High-level Agents
+Planning agent: The planning agent plays an important role in the multi-agent pipeline for managing problem sub-steps and determining which Low-level agents to deploy at each step.
+As shown in Figure 2, given the column candidates, current indexes, total storage budget, and budget already used, the planning agent first evaluates the stop criteria.
+If these criteria are not met, it outputs an action for the current sub-step: ‘Selection’, ‘Combination’, or ‘Revision’, each invoking a corresponding low-level agent.
+When the planning agent determines that the index selection is optimized or that no additional columns can be chosen within the budget, it triggers the ‘Stop’ action to conclude the optimization.
+Reflection agent: In practical use, it has been observed that planning agent’s strategies for low-level agents can be inefficient at times, as it may suffer from hallucinations or become trapped in local optima.
 For instance, the planning agent may repeatedly revise the current index recommendation, invoke the combination agent when no viable combinations exist, or produce invalid actions due to hallucinations.
-To enhance the planning process, a reflection agent is introduced to evaluate the current path, to identify and correct errors.
+To enhance the planning process, a reflection agent is introduced to evaluate the current path, with the goal of identifying and correcting errors.
 This helps the planning agent escape local optima and supports more efficient sub-step planning.
@@
 Example of possible suggestions can also be found in Figure 3(a).
-6.3. Multi-Agent Pipeline
-Once the global and local agents are developed, they can be integrated into a hierarchical multi-agent pipeline, enabling zero-shot scalable index recommendation.
+6.3. Multi-agent Planning and Coordination
+Once the high-level and low-level agents are developed, they can be integrated into a hierarchical multi-agent pipeline enabling efficient and effective zero-shot index recommendation.
 As shown in Algorithm 1, the multi-agent pipeline receives the storage budget and the column candidates as inputs.
 To balance efficiency with effectiveness, we introduce a hyperparameter , which defines the maximum number of planning sub-steps.
-At the beginning of each sub-step, the planning agent determines the next action.
-If additional indexes can be selected, the selection agent chooses a single-column index from the effectively represented column candidates.
-When combining existing indexes into a composite index could further improve workload performance, the combination agent refines the current recommendation by composing these indexes.
-Since the selected indexes may contain duplicates, regressions, or low-impact choices, the revision agent ensures effectiveness by removing such cases.
-After each action, the reflection agent reviews prior decisions and guides subsequent planning.
-Through the planning agent’s reasoning, if the index configuration is deemed optimal, iterations terminate early, and the final selection is output as the index recommendation.
-7. Experiments
-In this section, we will evaluate our multi-agent index advisor’s performance in terms of effectiveness, efficiency, and zero-shot generalisation ability.
-7.1. Experimental Setup
-7.1.1. Dataset.
-We select four different datasets from both hypothetical and real-world settings for our evaluations.
-TPC-H: The TPC-H dataset is a widely used benchmarking dataset for measuring the performance of database management systems.
-We generate the data with a scale of 10 (around 10 GB).
-The TPC-H benchmark has in total of 22 different query templates.
-In order to simplify the analysis and make a fair comparison of the methods, we follow previous methods (Siddiqui et al., 2022; Schlosser et al., 2019) to remove the query templates 2, 17, 20 that contribute significantly to the total execution time of the workload or have little potential improvement from indexes.
-The test workload for TPC-H has in total of 19 queries randomly generated from the 19 query templates left.
-TPC-DS: The TPC-DS dataset is another widely used benchmark with a more complicated database schema and query templates.
-Data with a scale of 10 (around 10 GB) is generated.
-The TPC-DS benchmark has 99 different query templates, and for the same reason in the TPC-H benchmark, we remove the query templates 4, 6, 9, 10, 11, 14, 23, 24, 32, 35, 36, 39, 41, 70, 86, and 95.
-The test workload for TPC-DS consists of 83 queries generated from the 83 query templates left.
-DSB: The DSB benchmark (Ding et al., 2021) is derived from the TPC-DS benchmark, while enhancing the TPC-DS with complex data distribution and challenging yet semantically meaningful query templates.
-DSB has in total of 15 templates for single block queries, 22 templates for multi-block queries, and another 15 templates for SPJ queries, especially evaluating techniques with limited capabilities.
-In our experiments, we only utilized the general-purpose 37 query templates in the first two groups and constructed a workload of 37 queries randomly generated from the templates.
-IMDB (JOB workload): The IMDB dataset (Maas et al., 2011) is a collection of information about movies, TV shows, and actors.
-Join order benchmark (JOB (Leis et al., 2015)) is an OLAP benchmark querying the IMDB dataset and designed to assess the capabilities of a database management system (DBMS) in optimizing and executing complex real-world join queries.
-The JOB workload has 113 real-world queries.
-7.1.2. Experimental Environments.
-We choose the popular DBMS PostgreSQL 14.2 to evaluate all the queries (Pos, [n.d.]).
-We follow the official documentaries and websites (TPC, [n.d.]; Leis et al., 2015; dsb, [n.d.]) to create PostgreSQL databases for all the above datasets.
-All the experiments are conducted on an isolated server with two Intel(R) Xeon(R) Platinum 8469C CPUs featuring 46 cores and 96 threads, along with 460GB of RAM.
-7.1.3. Baseline Methods.
-Heuristic-based baselines: Heuristic-based index advisors can be broadly categorized into two types by their optimization approaches.
-Those following an ADD strategy incrementally build a set of indexes by selecting candidates one by one.
-In our experiments, we include Anytime (Chaudhuri and Narasayya, 2020), DB2Advis (Skelley, 2000), Extend (Schlosser et al., 2019), and Relaxation (Bruno and Chaudhuri, 2005) as representatives of this group.
-On the other hand, AutoAdmin (Chaudhuri and Narasayya, 1997) and Drop (Choenni et al., 1993) serve as baselines for the DROP strategy, where advisors start with an initial set of indexes and iteratively eliminate those deemed inefficient.
-Learning-based baselines: We adopt two RL-based index advisors as our baselines, namely SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b).
-While SWIRL includes vast amounts of workloads during training to largely shorten the inference time, BALANCE utilizes transfer RL to increase the generalisability to the dynamic workload scenarios.
-Prompt-based baselines: The recent work LLMIdxAdvis (Zhao et al., 2025) proposes to recommend indexes by prompting the LLM with prepared demonstrations.
-Selecting the best demonstration for the input workload and calling the LLM helps LLMIdxAdvis achieve higher efficiency and generalisability.
-Due to the high computation cost for preparing new demonstrations for our workloads, we directly use LLMIdxAdvis’ prepared demonstrations since they cover all the databases in our experiment.
-7.1.4. Evaluation metrics.
-We adopt two metrics for our evaluations.
-The benefit-to-cost ratio metric is first introduced by (Skelley, 2000) and evaluates the absolute workload execution time reduction per unit storage used.
-The larger the benefit-to-cost ratio, the better an index advisor performs in both workload cost reduction and efficient storage use.
-The relative workload improvement is defined as the percentage of the workload’s actual execution time decrease before and after the construction of the recommended indexes.
-We measure all the workload execution time in seconds and storage cost in MegaBytes (MB).
-7.1.5. LLM and Training Setting.
-We leverage the GPT-4o-0806 (Achiam and et al., 2023) version within the OpenAI API as the default LLM backbone for the agents.
-Furthermore, we also assess our system’s generalizability across other LLMs, as detailed in Section 8.3.
-For the index indicator, we train the model on a server with one Nvidia A100 GPU.
+At the beginning of each sub-step, the planning agent determines the next action in line 7.
+If additional indexes can be selected, the selection agent chooses a single-column index from the effectively represented column candidates from line 9 to line 11.
+When combining existing indexes into a composite index could further improve workload performance, the combination agent in line 13 refines the current recommendation by composing these indexes.
+Since the selected indexes may contain duplicates, regressions, or low-impact choices, the revision agent ensures effectiveness by removing such cases once called in line 16.
+After each action, the reflection agent reviews prior decisions and provides guidance for subsequent planning.
+Through the planning agent’s reasoning in line 22, if the index configuration is deemed optimal, iterations terminate early, and the final selection is output as the index recommendation.
 | Dataset | TPC-H | TPC-DS | DSB | JOB | ||||||||||||
@@
 | LLMIdxAdvis | -0.013 | -0.025 | 0.684 | 0.203 | 2.136 | 0.754 | 1.642 | 0.394 | 0.061 | -0.630 | 0.246 | 0.250 | 0.943 | 0.202 | 0.126 | 0.090 |
-| MAAdivsor | 0.220 | 0.587 | 0.587 | 0.299 | 9.726 | 4.756 | 1.594 | 1.594 | 0.919 | 1.224 | 0.598 | 0.344 | 0.617 | 0.452 | 0.125 | 0.094 |
+| AMAZe | 0.220 | 0.587 | 0.587 | 0.299 | 9.726 | 4.756 | 1.594 | 1.594 | 0.919 | 1.224 | 0.598 | 0.344 | 0.617 | 0.452 | 0.125 | 0.094 |
+7. Experiment
+In this section, we will evaluate our multi-agent index advisor’s performance in terms of effectiveness, efficiency, and zero-shot generalisation ability.
+7.1. Experimental Setup
+Dataset. We select four different datasets from both hypothetical and real-world settings for our evaluations.
+TPC-H: We generate the TPC-H data with a scale of 10 (around 10 GB).
+The TPC-H benchmark has in total 22 different query templates.
+In order to simplify the analysis and make a fair comparison of the methods, we follow previous methods (Siddiqui et al., 2022b; Schlosser et al., 2019) to remove the query templates 2, 17, 20 that contribute significantly to the total execution time of the workload or have little potential improvement from indexes.
+The test workload for TPC-H has in total of 19 queries randomly generated from the 19 query templates left.
+TPC-DS: TPC-DS data with a scale of 10 (around 10 GB) is generated.
+The TPC-DS benchmark has 99 different query templates, and for the same reason in the TPC-H benchmark, we remove the query templates 4, 6, 9, 10, 11, 14, 23, 24, 32, 35, 36, 39, 41, 70, 86, and 95.
+The test workload for TPC-DS consists of 83 queries generated from the 83 query templates left.
+DSB: The DSB benchmark (Ding et al., 2021) is derived from the TPC-DS benchmark, while enhanced with complex data distribution and challenging query templates.
+DSB has in total 15 templates for single block queries, 22 templates for multi-block queries, and another 15 templates for SPJ queries, especially evaluating techniques with limited capabilities.
+In our experiments, we only utilized the general-purpose 37 query templates in the first two groups and constructed a workload of 37 queries randomly generated from the templates.
+IMDB (JOB workload): Join order benchmark (JOB (Leis et al., 2015)) is an OLAP benchmark querying the IMDB (Maas et al., 2011) dataset and designed to assess the capabilities of a database management system (DBMS) in optimizing and executing complex real-world join queries.
+The JOB workload has 113 real-world queries.
+Experiment Environments.
+We choose the popular DBMS PostgreSQL 14.2 to evaluate all the queries (Pos, [n.d.]).
+We follow the official documentaries and websites (TPC, [n.d.]; Leis et al., 2015; dsb, [n.d.]) to create PostgreSQL databases for all the above datasets.
+All the experiments are conducted on an isolated server with two Intel(R) Xeon(R) Platinum 8469C CPUs featuring 46 cores and 96 threads, along with 460GB of RAM.
+Baseline Methods.
+(1) Heuristic-based baselines: Heuristic-based index advisors can be broadly categorized into two types by their optimization approaches.
+Those following an ADD strategy incrementally build a set of indexes by selecting candidates one by one.
+In our experiments, we include Anytime (Chaudhuri and Narasayya, 2020), DB2Advis (Skelley, 2000), Extend (Schlosser et al., 2019), and Relaxation (Bruno and Chaudhuri, 2005) as representatives of this group.
+On the other hand, AutoAdmin (Chaudhuri and Narasayya, 1997) and Drop (Choenni et al., 1993) serve as baselines for the DROP strategy, where advisors start with an initial set of indexes and iteratively eliminate those deemed inefficient.
+We reproduce all the heuristic-based baselines based on the evaluation work (Kossmann et al., 2020).
+(2) Learning-based baselines: We adopt two RL-based index advisors as our baselines, namely SWIRL (Kossmann et al., 2022) and BALANCE (Wang et al., 2024b).
+While SWIRL includes vast amounts of workloads during training to largely shorten the inference time, BALANCE utilizes transfer RL to increase the generalisability to the dynamic workload scenarios.
+(3) Prompt-based baselines: The recent work LLMIdxAdvis (Zhao et al., 2025) proposes to recommend indexes by prompting the LLM with prepared demonstrations.
+Selecting the best demonstration for the input workload and calling the LLM helps LLMIdxAdvis achieve higher efficiency and generalisability.
+Due to the high computation cost for preparing new demonstrations for our workloads, we directly use LLMIdxAdvis’ prepared demonstrations since they cover all the databases in our experiment.
+Evaluation metrics.
+We adopt two metrics for our evaluations.
+The benefit-to-cost ratio metric is first introduced by (Skelley, 2000) and evaluates the absolute workload execution time reduction per unit storage used.
+The larger the benefit-to-cost ratio, the better an index advisor performs in both workload cost reduction and efficient storage use.
+The relative workload improvement is defined as the percentage of the workload’s actual execution time decrease before and after the construction of the recommended indexes.
+We measure all the workload execution time in seconds and storage cost in MegaBytes (MB).
+LLM and Training Setting.
+We leverage the GPT-4o-0806 (Achiam and et al., 2023) version within the OpenAI API as the default LLM backbone for the agents.
+Furthermore, we also assess our system’s generalizability across other LLMs, as detailed in Section 7.5.
+For the index indicator, we train the model on a server with one Nvidia A100 GPU.
 results_figure
-7.2. Effectiveness and Generalisation Evaluation
-As shown in Table 3, we compare our proposed MAAdvisor’s benefit-to-cost ratio with the baselines on four benchmarks.
+7.2. Generalization and Effectiveness Evaluation
+As shown in Table 3, we compare our proposed AMAZe’s benefit-to-cost ratio with the baselines on four benchmarks.
 All the methods are evaluated with four different storage budgets, taking 2%, 10%, 20%, and 50% of the database size.
-Our result reveals several key observations:
-(1) MAAdvisor outperforms all the baselines in most of the situations.
-The benefit-to-cost ratio for our method is the highest in 10 out of the 16 experiment settings, and remains the top tier for the rest 6 settings.
-Specifically, MAAdvisor doubles the index scores from the SOTA performance given 2%, 10%, and 50% storage budgets in the most complicated DSB benchmark.
-In addition, similar improvement can be observed in the 10% and 50% budget settings for the real-world benchmark JOB.
-It demonstrates that MAAdvisor can generalize to a variety of database schemas and workloads while maintaining excellent performance.
+Our results reveal several key observations:
+(1) AMAZe outperforms all the baselines in most of the situations.
+The benefit-to-cost ratio for our method is the highest in 10 out of the 16 experimental settings, and remains the top tier for the rest 6 settings, especially compared to the heuristic-based ßand prompt-based methods, which are also zero-shot.
+Specifically, AMAZe doubles the ratios from the SOTA performance given 2%, 10%, and 50% storage budgets in the most complicated DSB benchmark.
+Similar improvement can be observed in the 10% and 50% budget settings for the real-world benchmark JOB.
+This demonstrates AMAZe’s strong generalization capability to a variety of database schemas and workloads in zero-shot settings.
+(2) AMAZe successfully alleviates the index regression problem, which commonly occurs in the zero-shot heuristic-based methods.
+For example, Anytime suffers from the index regression in TPC-H and DSB under 10% budget, and similar problems can be observed for Extend in TPC-H and for SWIRL in TPC-DS under 10% budget.
+Our detailed analysis reveals that heuristic-based methods highly rely on the benefit estimations from optimizers, index regressions always happen when the index advisors have a limited storage budget and select the regressive indexes.
+For example, DB2Advis recommends indexes for TPC-H that lead to a nearly 50% execution time regression on query template 9 from 14.76 seconds to 26.60 seconds.
+Among the 13 indexable columns in query 9, only two columns appear in DB2Advis’ index recommendation.
+In addition, the composite indexes DB2Advis selects ‘I(C supplier.s_nationkey,C supplier.s_suppkey)’ and ‘I(C part.p_type,C part.p_partkey)’ are not efficient for JOIN conditions ‘ps_suppkey = l_suppkey’ and ‘p_partkey = l_partkey’.
+Unfortunately, the improvements for the indexes on other queries are not significant and therefore result in an overall regression.
+In contrast, our AMAZe mitigates the problem with the help of the specifically designed workload representation technique and low-level revision agent.
+(3) AMAZe mitigates the training instability or hallucination problem faced by learning-based / prompt-based methods.
+Our results show that the performance for the learning-based index advisors suffers from higher fluctuation across benchmarks.
+As an example, while BALANCE performs well on TPC-DS and DSB, it fails to beat heuristic-based baselines in the rest of the benchmarks due to the highly variable training process (Zhou et al., 2024).
+Similar variation happens in the prompt-based methods due to the hallucination of LLMs.
+As an example, LLMIdxAdvis’ performance on TPC-DS 20% budget is among the top tier, but the performance on the other budget settings is low.
+In contrast, our hierarchical multi-agent pipeline verifies and corrects the potential hallucinations during the index recommendation, effectively mitigating the problem.
+In order to further analyze the capability of our method, we plot the two metrics we used to compute the benefit-to-cost ratio, namely the relative workload improvement and the storage cost.
+From Figure 5, we can observe that:
+(1) AMAZe achieves the highest relative workload improvement in 8 out of in total 16 experiment settings.
+Given smaller storage budgets from 2% to 20%, AMAZe tends to recommend indexes that lead to the most workload relative reduction.
+However, for a larger budget like 50%.
+Our method prefers indexes with comparable improvement but much lower storage cost.
+It is also observed that in the experiment settings where our method did not achieve the highest benefit-to-cost ratio, for example, TPC-H 20% budget and JOB 20% budget, AMAZe can still recommend indexes that result in the largest workload execution time decrease, with a trade-off of higher storage cost.
+(2) AMAZe achieves SOTA performance even for small storage budgets.
+In contrast, heuristic-based baselines following the DROP heuristic fail to perform well for small storage budgets.
+The main reason is that the small-storage indexes are generally not as beneficial as large-storage indexes and are removed from the recommendation in early DROP stages.
+Our AMAZe assembles a hybrid of both ADD and DROP heuristics that iteratively selects and revises the index recommendation by multi-agent coordinations.
+(3) AMAZe with the multi-agent design outperforms the prompt-based baseline, especially in the most complicated DSB.
+The prompt-based LLMIdxAdvis performs well on TPC-H and JOB benchmarks but has a lower performance on complicated workloads and database schemas, like TPC-DS and DSB.
+AMAZe’s better performance comes from the efficient workload representation technique which summarizes the information from workloads and database schemas.
+In addition, the multi-agent pipeline decomposes the problem and simplifies the task for each LLM agent.
+(4) AMAZe better controls the storage budget usage during the recommendation.
+It is observed that LLMIdxAdvis struggles to control the storage budget used in TPC-DS and DSB.
+As an example in TPC-DS, LLMIdxAdvis uses 1356.38 MB under 10% budget, which exceeds the budget, and a lower 1122.97 MB under 20% budget.
+The main reason is that although the available storage budget is an input in LLMIdxAdvis, LLMs are not given the storage cost when selecting new indexes.
+In contrast, AMAZe’s planning and selection of indexes is always storage-aware since the estimated storage cost is recorded in each one of the column candidates.
 | Dataset | TPC-H | TPC-DS | DSB | JOB | ||||||||||||
@@
 | LLMIdxAdvis | 210.23 | 148.95 | 216.63 | 203.83 | 153.35 | 167.76 | 412.48 | 138.95 | 134.77 | 123.45 | 146.10 | 177.62 | 141.86 | 165.49 | 170.04 | 174.96 |
-| MAAdivsor | 31.85 | 44.31 | 122.21 | 156.82 | 51.00 | 85.75 | 106.71 | 133.21 | 43.23 | 83.67 | 115.38 | 178.82 | 69.22 | 66.18 | 106.62 | 211.56 |
-(2) Index regressions happen for the zero-shot heuristic-based methods, for example, Anytime in TPC-H and DSB under 10% budget.
-Similar results can be observed for Extend in TPC-H and for SWIRL in TPC-DS under 10% budget.
-Since heuristic-based methods highly rely on the benefit estimations from optimizers, index regressions always happen when the index advisors have a limited storage budget and select the regressive indexes.
-As an example, DB2Advis recommends indexes for TPC-H that lead to a nearly 50% execution time regression on query template 9 from 14.76 seconds to 26.60 seconds.
-Among the 13 indexable columns in query 9, only two columns appear in DB2Advis’ index recommendation.
-In addition, the composite indexes DB2Advis selects ‘I(C supplier.s_nationkey,C supplier.s_suppkey)’ and ‘I(C part.p_type,C part.p_partkey)’ are not efficient for JOIN conditions ‘ps_suppkey = l_suppkey’ and ‘p_partkey = l_partkey’.
-Unfortunately, the improvements for the indexes on other queries are not significant and therefore result in an overall regression.
-In contrast, MAAdvisor successfully alleviates the index regression problem while maintaining the zero-shot generalisation ability, with the help of the specifically designed workload representation technique and local revision agent.
-(3) Aligned with (Zhou et al., 2024), it is observed that the performance for the learning-based index advisors suffers from higher fluctuation across benchmarks compared to heuristic-based methods.
-As an example, while BALANCE performs well on TPC-DS and DSB, it fails to beat heuristic-based baselines for the rest of the benchmarks.
-The main reason is that the training process for the learning-based method is highly variable.
-The hallucination of LLMs is also a similar issue that may affect the variance of the performance.
-As an example, LLMIdxAdvis’ performance on TPC-DS 20% budget is among the top tier, but the performance on the other budget settings is low.
-In contrast, with the help of the hierarchical multi-agent pipeline, our zero-shot MAAdvisor maintains a higher performance across the datasets without training and is robust against hallucination.
-In order to further analyze the capability of our method, we plot the two metrics we used to compute the benefit-to-cost ratio, namely the relative workload improvement and the storage cost.
-From Figure 5, we can observe that:
-| Avg. Tokens (Cost $) | TPC-H | TPC-DS | DSB | JOB |
-| MAAdvisor | 12165(0.03) | 68035(0.17) | 66926(0.17) | 33117(0.08) |
-| LLMIdxAdvis | 38736(0.10) | 38859(0.10) | 37361(0.10) | 35619(0.09) |
-(1) MAAdvisor achieves the highest relative workload improvement in 8 out of in total 16 experiment settings.
-Given smaller storage budgets from 2% to 20%, MAAdvisor tends to recommend indexes that lead to the most workload relative reduction.
-However, for a larger budget like 50%.
-Our method prefers indexes with comparable improvement but much lower storage cost.
-It is also observed that in the experiment settings where our method did not achieve the highest benefit-to-cost ratio, for example, TPC-H 20% budget and JOB 20% budget, MAAdvisor can still recommend indexes that result in the largest workload execution time decrease, with a trade-off of higher storage cost.
-(2) Heuristic-based baselines following the DROP heuristic fail to perform well for small storage budgets.
-The main reason is that the small-storage indexes are generally not as beneficial as large-storage indexes and are removed from the recommendation in early DROP stages.
-Our MAAdvisor assembles the ADD heuristic that iteratively selects indexes into the recommendation by calling the selection agent.
-Therefore, MAAdvisor can achieve SOTA performance even for small storage budgets.
-(3) The prompt-based method LLMIdxAdvis performs well on TPC-H and JOB benchmarks but has a lower performance on TPC-DS and DSB.
-As also specified in (Zhao et al., 2025), directly calling LLMs to select indexes fails to excel in these complicated workloads and database schemas.
-In contrast, MAAdvisor outperforms all the baselines, especially in DSB, with our special design.
-The workload representation technique efficiently summarizes the information from workloads and database schemas.
-In addition, the multi-agent pipeline decomposes the problem into sub-steps handled by specialized agents, simplifying the task for each LLM agent to solve.
-(4) We find LLMIdxAdvis also struggles to control the storage budget used in TPC-DS and DSB.
-In TPC-DS, LLMIdxAdvis uses 1356.38 MB under 10% budget, which exceeds the budget, and a lower 1122.97 MB under 20% budget.
-The main reason is that although the available storage budget is an input for LLM in LLMIdxAdvis, the inputs do not specify the storage cost for selecting new indexes.
-In contrast, MAAdvisor’s planning and selection of indexes is always storage-aware since the estimated storage cost is recorded in each one of the column candidates.
-queryid
+| AMAZe | 31.85 | 44.31 | 122.21 | 156.82 | 51.00 | 85.75 | 106.71 | 133.21 | 43.23 | 83.67 | 115.38 | 178.82 | 69.22 | 66.18 | 106.62 | 211.56 |
 7.3. Efficiency Evaluation
-In Table 4, we record the algorithm inference runtime for all the methods to evaluate their efficiency.
-Through analyzing the inference runtime of different methods, we have the following key observations:
-(1) We can observe that our MAAdvisor’s algorithm runtime outperforms the prompt-based baseline LLMIdxAdvis in most of the settings.
-The reason is that with the multi-agent pipeline, our method’s only time overheads come from invoking the LLM API and refreshing column candidates, whereas LLMIdxAdvis faces an efficiency bottleneck due to the added cost of identifying the best demonstrations for each selection.
+Table 4 reports the algorithm inference runtime for all the methods to evaluate their efficiency. Through analyzing the inference runtime of different methods, we make the following observations:
+(1) AMAZe’s runtime outperforms the prompt-based baseline LLMIdxAdvis in most of the settings.
+The reason is that with the multi-agent pipeline, our method’s main computational overhead comes from invoking the LLM API and updating column candidates, whereas LLMIdxAdvis faces an efficiency bottleneck due to the added cost of identifying the best demonstrations for each selection.
 In addition, LLMIdxAdvis suffers from high latency during its demonstration construction stage.
-It is reported that 17 hours and $90 are taken to collect the demonstrations (Zhao et al., 2025).
-It is also important to note that LLMIdxAdvis’ time cost, given increasing storage budgets, does not change a lot since they call the same demonstration match algorithm.
-In contrast, our MAAdvisor’s algorithm runtime may increase as the budget increases.
+17 hours and $90 are reported to be taken to collect the demonstrations (Zhao et al., 2025).
+It is also note-worthy that LLMIdxAdvis’ time cost, given increasing storage budgets, does not change a lot since they call the same demonstration match algorithm.
+In contrast, AMAZe’s runtime may increase as the budget increases.
 It is because our index recommendation is storage-aware.
-The more budget given, the more local agents are planned by the planning agent.
-However, MAAdvisor’s time cost under the largest 50% budget is still lower than LLMIdxAdvis.
+The more budget given, the more low-level agents are planned by the planning agent.
+However, AMAZe’s time cost under the largest 50% budget is still lower than LLMIdxAdvis.
 (2) The learning-based methods have significantly faster inference time compared to the other methods.
@@
 As an example, SWIRL’s training times for the TPC-H, TPC-DS, and JOB datasets are 1226.136, 3373.32, and 3635.00 seconds, respectively.
-In addition, both learned methods require an initial training set of data collected from the database, which is difficult to acquire in a zero-shot.
-(3) Heuristic-based advisors on larger workloads generally suffer from much higher latency.
-The algorithm runtime for Relaxation on TPC-DS with an 83-query workload is more than 100 times the runtime on the TPC-H benchmark with 19 queries.
-For a larger workload and more complicated database, MAAdvisor presents a significant efficiency advantage.
-Although the DB2Advis baseline demonstrates a shorter runtime compared to our MAAdvisor, it trades off with a lower workload performance by using a simpler greedy optimization.
-pipeline_example
+In addition, both learned methods require an initial training set of data collected from the database, which is difficult to acquire in a zero-shot setting.
+(3) AMAZe presents a significant efficiency advantage for larger workloads and more complicated database schemas.
+Heuristic-based advisors on larger workloads generally suffer from much higher latency.
+As an example, the algorithm runtime for Relaxation on TPC-DS with 83 queries’ workload is more than 100 times the runtime on the TPC-H benchmark with 19 queries.
+Although the DB2Advis baseline shows a shorter runtime compared to our AMAZe, it trades off with a lower workload performance by using a simpler greedy optimization.
 (4) It is important to note that for the methods AutoAdmin and Drop following the DROP heuristic, the hyperparameter constraint for these methods is designed to be the maximum number of indexes to recommend.
@@
 As an example, the total time of recommending indexes for TPC-DS under a 10% budget is 689.05 seconds for AutoAdmin and 5334.80 seconds for Drop.
-We further evaluate the efficiency of our MAAdvisor with the LLM baseline LLMIdxAdvis in terms of monetary cost.
-The average number of input tokens and corresponding API cost for MAAdvisor and LLMIdxAdvis is also recorded in Table 5.
-It is observed that LLMIdxAdvis’ input lengths are more stable than MAAdvisor across benchmarks.
+We further evaluate the performance of AMAZe with the LLM baseline LLMIdxAdvis in terms of monetary cost.
+The average number of input tokens and the corresponding API cost for AMAZe and LLMIdxAdvis are recorded in Table 5.
+It is observed that LLMIdxAdvis’ input lengths are more stable across benchmarks.
 The main reason is that LLMIdxAdvis uses a similar number of demonstrations in all benchmarks.
-In contrast, MAAdvisor’s input length depends on database complication and the number of column candidates.
-As an example, MAAdvisor takes 4 times more tokens for TPC-DS with 87 column candidates than TPC-H with 15 column candidates.
-Considering the complexity of the index recommendation task, MAAdvisor’s additional cost on more complicated workloads is totally acceptable.
+In contrast, AMAZe’s input length depends on database complication and the number of column candidates.
+As an example, AMAZe takes 4 times more tokens for TPC-DS with 87 column candidates than TPC-H with 15 column candidates.
+Considering the complexity of index recommendation task, AMAZe’s higher cost on more complicated workloads is acceptable.
+| Avg. Tokens (Cost $) | TPC-H | TPC-DS | DSB | JOB |
+| AMAZe | 12165(0.03) | 68035(0.17) | 66926(0.17) | 33117(0.08) |
+| LLMIdxAdvis | 38736(0.10) | 38859(0.10) | 37361(0.10) | 35619(0.09) |
+pipeline_example
+queryid
 7.4. Qualitative Evaluation
-An example of MAAdvisor’s planning steps and index recommendation on TPC-H 20% budget can be found in Figure 7.
-We can observe that MAAdvisor efficiently recommends the indexes within five sub-steps, including selection and combination.
-Since the index selected is just within the storage budget, the revision agent is not called by the planning agent.
-It is also important to note that MAAdvisor adopts the same storage cost estimation as the other baselines to increase efficiency.
+An example of AMAZe’s planning steps and index recommendation on TPC-H 20% budget can be found in Figure 6.
+We observe that AMAZe efficiently recommends the indexes within five sub-steps, including selection and combination.
+Since the selected index is within the storage budget, the revision agent is not called by the planning agent.
+Note that AMAZe adopts the same storage cost estimation as the other baselines.
 Therefore, the actual storage cost of the output is smaller than the estimation in the planning phase.
-MAAdvisor can also opt to construct and calculate the actual index storage cost at each sub-step at a cost of a higher algorithm runtime.
-In order to further analyze how well MAAdvisor addresses the index regression problem, we pick out the queries in each benchmark having any methods’ recommended indexes under 20% budget leading to a more than one-second’s regression.
+AMAZe can also opt to construct and calculate the actual index storage cost at each sub-step at a cost of a higher algorithm runtime.
+In order to further analyze how well AMAZe addresses the index regression problem, we pick out the queries in each benchmark having any methods’ recommended indexes under 20% budget leading to a more than one second’s regression.
 As shown in Figure 7, it is observed that DB2Advis has the most regression cases, trading off the high algorithm efficiency.
-Most heuristic-based advisors suffer from regressions on the most complicated DSB benchmark, since their optimization relies solely on the optimizer’s estimations.
+Most heuristic-based advisors suffer from regressions on the most complicated DSB benchmark, since their optimization relies solely on the optimizer estimations.
 It is important to note that, at the workload level, significant improvements on some queries with a cost of regression on some other queries is acceptable as long as the overall execution time is decreased.
-On the other hand, although MAAdvisor does not result in the highest query time improvement on all mentioned queries, the revision agent manages to avoid significant regressions and therefore leads to a higher workload-level performance.
-8. Ablation Studies
-In this section, we first conduct an ablation study to evaluate the effectiveness of MAAdvisor’s workload representation technique and multi-agent pipeline.
-We will then analyze MAAdvisor’s generalisation ability over the LLM backbone and workload scale.
-8.1. Effectiveness of Workload Representation
-To validate the effectiveness of MAAdvisor’s workload representation step, we perform an ablation study in which the database schema and raw workload queries are directly input to the LLM agents.
+On the other hand, although AMAZe does not result in the highest query time improvement on all mentioned queries, the revision agent manages to avoid significant regressions and therefore leads to a higher workload-level performance.
+7.5. Ablation Studies
+In this section, we first conduct an ablation study to evaluate the effectiveness of AMAZe’s workload representation technique and multi-agent pipeline.
+We will then analyze AMAZe’s generalisation ability over the LLM backbone and workload scale.
+Effectiveness of Workload Representation.
+To validate the effectiveness of AMAZe’s workload representation step, we perform an ablation study in which the database schema and raw workload queries are directly input to the LLM agents.
 Instead of selecting from column candidates, the selection agent is now prompted to select one possible column index from the SQL queries.
@@
 In addition, the input context for the agents will be much longer, and LLMs’ hallucination and misunderstanding will increase significantly.
-| Avg. Relative workload improvement | (%) |
-| MAAdvisor (TPC-H 1x) | 61.33 |
-| - Raw Workload | 29.52 |
-| - One-call Recommendation | 11.20 |
-| - Raw Workload and One-call Recommendation | 10.58 |
-| - w/o Regression Indicator | 58.21 |
-| MAAdvisor (GPT-3.5-turbo) | 60.01 |
-| MAAdvisor (QWen3 32b) | 60.34 |
-| MAAdvisor (QWen3 14b) | 40.86 |
-| MAAdvisor (DeepCoder 14b) | 40.27 |
-| MAAdvisor (TPC-H 20x) | 61.31 |
-8.2. Effectiveness of Multi-agent pipeline
+Effectiveness of Multi-agent pipeline.
 To explore the effectiveness of our multi-agent pipeline, we prompt the LLM to directly select from the column candidates and construct indexes in one step.
@@
 The relative workload improvement further drops to 10.58%, showing that both our workload representation technique and the multi-agent pipeline are important in solving the index recommendation problem.
-8.3. Effectiveness of Regression Indicator
+Effectiveness of Regression Indicator.
 We further test the effectiveness of the specially trained regression indicator embedded in the revision agent.
-We apply the MAAdvisor to the TPC-H benchmark with the regression indicator removed.
+We apply the AMAZe to the TPC-H benchmark with the regression indicator removed.
 The revision agent will then only rely on the provided expert experience to revise the current index recommendation.
-We can observe that MAAdvisor’s performance only suffers from a small drop.
+We can observe that AMAZe’s performance only suffers from a small drop.
 The main reason is that for smaller budgets, the index recommendation with a multi-agent pipeline is straightforward without too many revisions required, as shown in Figure 6.
 In addition, the revision agent alone can handle most of the severely regressive indexes by analysing the index utility estimation and learning from the expert experiences.
-8.4. Different LLM Backbones
-To evaluate MAAdvisor’s performance given different LLM backbones, we adopt one LLM API GPT-3.5-turbo (Brown et al., 2020), and three open-source LLMs, namely Qwen3-32B (Yang et al., 2025), Qwen3-14B, and DeepCoder-14B (Luo et al., 2025).
-DeepCoder-14B is the latest LLM that is specially finetuned on code-related data and achieves SOTA performance in code-related tasks.
+| Avg. Relative workload improvement | (%) |
+| AMAZe (TPC-H 1x) | 61.33 |
+| - Raw Workload | 29.52 |
+| - One-call Recommendation | 11.20 |
+| - Raw Workload and One-call Recommendation | 10.58 |
+| - w/o Regression Indicator | 58.21 |
+| AMAZe (GPT-3.5-turbo) | 60.01 |
+| AMAZe (QWen3 32b) | 60.34 |
+| AMAZe (QWen3 14b) | 40.86 |
+| AMAZe (DeepCoder 14b) | 40.27 |
+| AMAZe (TPC-H 20x) | 61.31 |
+Different LLM Backbones To evaluate AMAZe’s performance given different LLM backbones, we adopt an API GPT-3.5-turbo (Brown et al., 2020), and three open-source LLMs, namely Qwen3-32b (Yang et al., 2025), Qwen3-14b and DeepCoder-14b (Luo et al., 2025).
+DeepCoder 14b is the latest LLM that is specially finetuned on code-related data and achieves SOTA performance in code-related tasks.
 The results of adopting different LLM backbones are shown in Table 6.
-Since MAAdvisor relies on the reasoning ability of the LLM, we can observe that GPT-3.5-turbo and Qwen3-32B have comparable performance to GPT-4o-0806 and outperform Qwen3-14B.
-Smaller LLMs can also perform index recommendation following our MAAdvisor pipeline, but with a lower performance.
-It is important to note that the DeepCoder, specially fine-tuned on code-related data, underperforms the Qwen3 model with the same parameter size, showing that the coding ability of LLMs is not a must, given the rich information provided by the column candidates.
-In contrast, the reasoning ability of the LLMs is more important.
-8.5. Different Workload Scales
-In real-world applications, the workloads may always be updating, and such a situation is usually referred to as a dynamic setting.
+Since AMAZe relies on the reasoning ability of the LLM, we observe that GPT-3.5-turbo and Qwen3-32b have performance comparable to GPT-4o-0806 and outperform Qwen3-14b.
+Smaller LLMs can also perform index recommendation following AMAZe pipeline, but with a lower performance.
+It is note-worthy that the DeepCoder, specially fine-tuned on code-related data, underperforms the Qwen3 model with the same parameter size, showing that the coding ability of LLMs is not a must, given the rich information provided by the column candidates.
+In contrast, the reasoning ability of LLMs is more important.
+Different Workload Scales In real-world applications, the workloads may always be updating, and such a situation is usually referred to as a dynamic setting.
 In order to simulate the change of scales, we construct another TPC-H workload of 380 queries named TPC-H 20x, where 20 random queries are generated for each query template.
-As shown in Table 6, MAAdvisor’s performance is robust against different workload scales.
+As shown in Table 6, AMAZe’s performance is robust against different workload scales.
 In addition, since the number of column candidates for the TPC-H 20x is almost the same as TPC-H 1x, the algorithm runtime for TPC-H 20x only increases from an average of 88.80 seconds to 92.91 seconds.
 8. Conclusion and Discussion
-In this work, we present MAAdvisor, a novel zero-shot index advisor that leverages the reasoning power of large language models through a hierarchical multi-agent pipeline.
-By representing the workload using column candidates and decomposing the index recommendation task into coordinated sub-steps handled by specialized agents, MAAdvisor achieves superior performance, strong generalisability, and high efficiency without the need for database-specific training.
-Extensive experiments across diverse benchmarks demonstrate that MAAdvisor not only outperforms traditional heuristic and learning-based approaches but also surpasses existing prompt-based LLM methods.
+In this work, we present AMAZe, a novel zero-shot index advisor that leverages the reasoning power of large language models through a hierarchical multi-agent pipeline.
+By representing the workload using column candidates and decomposing the index recommendation task into coordinated sub-steps handled by specialized agents, AMAZe achieves superior performance, strong generalisability, and high efficiency without the need for database-specific training.
+Extensive experiments across diverse benchmarks demonstrate that AMAZe not only outperforms traditional heuristic and learning-based approaches but also surpasses existing prompt-based LLM methods.
 It is also important to acknowledge the current limitations and outline potential directions for future work.
-The efficiency of MAAdvisor may be further improved through the design of more precise and effective instruction sets.
+The efficiency of AMAZe may be further improved through the design of more precise and effective instruction sets.
 Additionally, fine-tuning the agent reasoning pipeline (Wang et al., 2024a) offers a promising avenue to reduce planning overhead and facilitate deployment on local LLMs.
@@
 https://arxiv.org/abs/2005.14165
+- Brucato et al.
+(2024) Matteo Brucato, Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri.
+2024.
+Wred: Workload Reduction for Scalable Index Tuning.
+Proc.
+ACM Manag.
+Data 2, 1, Article 50 (March 2024), 26 pages.
+https://doi.org/10.1145/3639305
 - Bruno and Chaudhuri (2005) Nicolas Bruno and Surajit Chaudhuri.
@@
 https://doi.org/10.1145/1066157.1066184
+- Chang et al.
+(2024) Zhuo Chang, Xinyi Zhang, Yang Li, Xupeng Miao, Yanzhao Qin, and Bin Cui.
+2024.
+MFIX: An Efficient and Reliable Index Advisor via Multi-Fidelity Bayesian Optimization.
+In 2024 IEEE 40th International Conference on Data Engineering (ICDE).
+4343–4356.
+https://doi.org/10.1109/ICDE60146.2024.00331
 - Chaudhuri and Narasayya (2020) Surajit Chaudhuri and Vivek Narasayya.
@@
 https://doi.org/10.1145/3299869.3324957
-- He et al.
-(2016) Ji He, Jianshu Chen, Xiaodong He, Jianfeng Gao, Lihong Li, Li Deng, and Mari Ostendorf.
-2016.
-Deep Reinforcement Learning with a Natural Language Action Space.
-In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Katrin Erk and Noah A.
-Smith (Eds.).
-Association for Computational Linguistics, Berlin, Germany, 1621–1630.
-https://doi.org/10.18653/v1/P16-1153
 - Jha et al.
@@
 https://doi.org/10.1145/3571730
+- Kossmann et al.
+(2020) Jan Kossmann, Stefan Halfpap, Marcel Jankrift, and Rainer Schlosser.
+2020.
+Magic mirror in my hand, which is the best in the land?
+an experimental evaluation of index selection algorithms.
+Proc.
+VLDB Endow.
+13, 12 (July 2020), 2382–2395.
+https://doi.org/10.14778/3407790.3407832
 - Kossmann et al.
@@
 - Perera et al.
-(2021) R.
+(2021a) R.
 Malinga Perera, Bastian Oetomo, Benjamin I.
@@
 Rubinstein, and Renata Borovica-Gajic.
-2021.
+2021a.
 DBA bandits: Self-driving index tuning under ad-hoc, analytical workloads with safety guarantees.
@@
 IEEE, 600–611.
+https://doi.org/10.1109/ICDE51399.2021.00058
+- Perera et al.
+(2021b) R.
+Malinga Perera, Bastian Oetomo, Benjamin I.
+P.
+Rubinstein, and Renata Borovica-Gajic.
+2021b.
+DBA bandits: Self-driving index tuning under ad-hoc, analytical workloads with safety guarantees.
+In 2021 IEEE 37th International Conference on Data Engineering (ICDE).
+600–611.
 https://doi.org/10.1109/ICDE51399.2021.00058
@@
 https://arxiv.org/abs/1707.06347
-- Shen et al.
-(2023) Yongliang Shen, Kaitao Song, Xu Tan, Dongsheng Li, Weiming Lu, and Yueting Zhuang.
-2023.
-HuggingGPT: Solving AI Tasks with ChatGPT and its Friends in Hugging Face.
-In Advances in Neural Information Processing Systems, A.
-Oh, T.
-Naumann, A.
-Globerson, K.
-Saenko, M.
-Hardt, and S.
-Levine (Eds.), Vol.
-36.
-Curran Associates, Inc., 38154–38180.
-https://proceedings.neurips.cc/paper_files/paper/2023/file/77c33e6a367922d003ff102ffb92b658-Paper-Conference.pdf
 - Shi et al.
@@
 - Siddiqui et al.
-(2022) Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri.
-2022.
+(2022a) Tarique Siddiqui, Saehan Jo, Wentao Wu, Chi Wang, Vivek Narasayya, and Surajit Chaudhuri.
+2022a.
+ISUM: Efficiently Compressing Large and Complex Workloads for Scalable Index Tuning.
+In Proceedings of the 2022 International Conference on Management of Data (Philadelphia, PA, USA) (SIGMOD ’22).
+Association for Computing Machinery, New York, NY, USA, 660–673.
+https://doi.org/10.1145/3514221.3526152
+- Siddiqui et al.
+(2022b) Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri.
+2022b.
 DISTILL: low-overhead data-driven techniques for filtering and costing indexes for scalable index tuning.
@@
 https://aclanthology.org/2024.findings-emnlp.459
+- Wang and Yoneki (2024) Taiyi Wang and Eiko Yoneki.
+2024.
+IA2: Leveraging Instance-Aware Index Advisor with Reinforcement Learning for Diverse Workloads.
+In Proceedings of the 4th Workshop on Machine Learning and Systems (Athens, Greece) (EuroMLSys ’24).
+Association for Computing Machinery, New York, NY, USA, 10–17.
+https://doi.org/10.1145/3642970.3655839
 - Wang et al.
@@
 https://arxiv.org/abs/2303.03846
+- Wu et al.
+(2022) Wentao Wu, Chi Wang, Tarique Siddiqui, Junxiong Wang, Vivek Narasayya, Surajit Chaudhuri, and Philip A.
+Bernstein.
+2022.
+Budget-aware Index Tuning with Reinforcement Learning.
+In Proceedings of the 2022 International Conference on Management of Data (Philadelphia, PA, USA) (SIGMOD ’22).
+Association for Computing Machinery, New York, NY, USA, 1528–1541.
+https://doi.org/10.1145/3514221.3526128
 - Yang et al.